### v1

# Energy, enstrophy and helicity transfers in polymeric turbulence

Abstract
We characterise the scale-by-scale transfers of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence using direct numerical simulations. The microscale Reynolds number is set to $Re_{\lambda}\approx 460$, and the Deborah number $De=\tau_{p}/\tau_{f}$ is varied between $1/9\leq De\leq 9$; $\tau_{p}$ is the polymeric relaxation time and $\tau_{f}$ is the turnover time of the largest scales of the flow. The study relies on the exact scale-by-scale budget equations (derived from the the governing model equations) for energy, enstrophy and helicity, which account for the back-reaction of the polymers on the flow. Polymers act as a sink/source in the flow, and provide alternative routes for the scale-by-scale transfers of the three quantities, whose relevance changes with $De$. We find that polymers deplete the nonlinear energy cascade mainly at smaller scales, by weakening both the extreme forward as well as reverse local events. The new polymer-driven energy flux dominates at small scales for $De\geq 1$, and on average transfers energy from larger to smaller scales with localised backscatter events. Polymers weaken the stretching of vorticity with the enstrophy being mainly generated by the fluid-polymer interaction, especially when $De\geq 1$. Accordingly, an inspection of the small-scale flow topology shows that polymers favour events with two-dimensional state of straining, and promote/inhibit extreme extension/rotation events: in polymeric turbulence shear and planar extensional flows are more probable. The helicity injected at the largest scales shows a similar transfer process to as energy, being mainly driven by the nonlinear cascade at large scales and by the polymer-driven flux at small scales. Polymers are found to favour events that break the small-scale mirror symmetry, with the relative helicity monotonically increasing with $De$ at all scales.
keywords:
1 Introduction
A small concentration of polymers in a turbulent flow results in a substantial decrease of drag, as is well documented in several experimental, numerical and theoretical works (Toms, 1948; Lumley, 1969; Berman, 1978; Benzi & Ching, 2018). In a turbulent wall-bounded flow, the polymeric chains influence the energy- momentum transfers in the real space and in the space of scales, giving rise to an increased buffer region that eventually leads to drag reduction (see for example Warwaruk & Ghaemi, 2024). It is known that large drag reduction may occur without a substantial reduction of the turbulent kinetic energy, which may indeed be either smaller or larger compared to the purely Newtonian counterpart (Tsinober, 1990). Polymeric turbulence, therefore, is not necessarily associated with the suppression of turbulence, but rather with qualitative/quantitative changes of its structure. Despite the large interest, however, the way polymeric additives modulate turbulence is not completely understood yet, and their influence on the statistical nature of the velocity fluctuations is still under debate. In this work, we consider the idealised set up of homogeneous and isotropic turbulence (HIT), and investigate the influence of polymeric additives on the cascades of energy, helicity and enstrophy in the space of scales.
1.1 Polymers in homogeneous and isotropic turbulence
Two dimensionless numbers are needed to characterise turbulent flows of dilute polymeric solutions in a triperiodic box. They are the Reynolds number $Re$, which estimates the importance of the inertial term compared to the viscous one in the Navier–Stokes equations, and the Deborah number $De$ that quantifies the unique polymer relaxation time-scale with respect to the large time scale of the flow. A large De therefore means the polymers are more elastic, and extensible. Turbulent drag reduction has been mainly observed for large Re and De, with an evident departure from the Kolmogorov predictions already in the idealised setting of HIT (Tabor & Gennes, 1986; Bhattacharjee & Thirumalai, 1991; Fouxon & Lebedev, 2003). At rather low Re of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 80$ ($u^{\prime}$ is the typical velocity fluctuation, $\lambda$ is the Taylor length-scale and $\nu$ is the kinematic fluid viscosity), Perlekar et al. (2010) found that in the presence of the polymers the energy content decreases at the intermediate scales and significantly increases at the smallest scales, in the deep dissipative range. They also observed that the fluid dissipation monotonically decreases as $De$ increases, although they fixed the injected energy using the forcing introduced by Lamorgese et al. (2005), hinting that polymers give origin to an alternative route to the classical energy cascade.
At large $Re$ and $De$, the polymer additives modify the nature of the energy cascade and result in a significant alteration of the energy distribution amongst scales. For large enough separation between the energy injection scale $\mathcal{L}$ and the Kolmogorov scale $\eta$, there exists an intermediate scale $r_{p}$ that separates two different regimes. For $r_{p}\leq r\leq\mathcal{L}$ the energy cascade resembles that of a Newtonian flow, with the classical second-order structure function scaling $S_{2}(r)\sim r^{2/3}$ predicted by the Kolmogorov theory. For $\eta\leq r\leq r_{p}$, the elastic range of scales, energy is transferred among scales by the polymeric microstructure as well, and the second-order structure function $S_{2}$ increases faster than $r^{2/3}$, deviating from the Kolmogorov predictions (De Angelis et al., 2005). Recently, Zhang et al. (2021) and Rosti et al. (2023) provided experimental and numerical evidence that, for large enough $De$, there is a range of scales $\eta\lesssim r\lesssim r_{p}$ where $S_{2}(r)\sim r^{\xi}$, with $\xi\approx 1.3\approx 4/3$. By studying the scale-by-scale energy budget in Fourier space, Rosti et al. (2023) observed that the energy flux at only the large scales is dominated by the advective nonlinearity for large enough $De$ (as for purely Newtonian turbulence), and that the polymeric stresses become dominant at smaller scales. Interestingly, they also observed that $r_{p}$ has a non-monotonic dependence on $De$, and follows the cross-over scale between the nonlinear flux and the non-Newtonian flux. As a consequence, the width of the elastic range of scales does not show a monotonic dependence on the $De$, and is maximum for $De\approx 1$. More recently, Singh & Rosti (2024) investigated how polymeric turbulence depends on $De$ and $Re$, linking the low-$Re$ regime described by Singh et al. (2024) with the large-$Re$ regime described by Rosti et al. (2023). They found that in polymeric turbulence the dissipation field is intermittent and shows a qualitative similar distribution as in purely Newtonian turbulence.
However, the influence of the polymers on the average picture of the energy cascade does not provide an exhaustive understanding of polymeric turbulence. On one side, indeed, already in Newtonian turbulence it is well known that the (average) forward energy cascade from larger to smaller scales is accompanied by local backscatter events where energy is transferred from smaller to larger scales. On the other side, despite energy is a central quantity in most of the theories, a complete understanding of polymeric turbulence requires to also focus on other small-scale related quantities such as enstrophy and helicity, which are strongly related to important dynamical processes such as vortex stretching and dissipation. In this work we do a step in this direction. We consider polymeric homogeneous and isotropic turbulence (PHIT) at large $Re$ and over a wide range of $De$, and, introducing a suitable framework, we address the following questions: How do polymers modulate the local direct/inverse energy transfers? How intermittent is the polymer-driven energy transfer route? Is it characterised by local backscatter events? How do polymers modulate the transfers of enstrophy and helicity among scales and the related dynamical processes?
1.2 Enstrophy and Helicity
When studying turbulent flows, the kinetic energy has always been of a great interest: it is an invariant of the inviscid Navier–Stokes equations and it is central to most of the theories developed for high Re turbulence (Frisch, 1995). However, it is not the only key quantity, and a general description of turbulence also requires the characterisation of quantities related with the smallest scales of the flow. In fact, it is well known that the energy cascade is statistically related to dissipation (Kolmogorov, 1941). Under homogeneity and isotropy, in the inertial range of scales, the nonlinear flux and the dissipation $\varepsilon_{f}$ are related by the celebrated $4/5$-th law, which directly descends from the Kármán-Howarth equation (Frisch, 1995; Pope, 2000). The smallest scales of the flow are characterised by large velocity gradients that lead to large values of the vorticity $\bm{\omega}=\bm{\nabla}\times\bm{u}$, that is central in the definition of other scalar quantities as enstrophy $\omega^{2}=\bm{\omega}\cdot\bm{\omega}$ and helicity $h=\bm{\omega}\cdot\bm{u}$. Although polymers are known to significantly alter the small scales (Perlekar et al., 2010) by suppressing events of large vorticity and strain (Liberzon et al., 2005, 2006), an exhaustive characterisation is missing. A complete understanding of how polymers modulate turbulence thus requires a deeper investigation of their influence on the enstrophy and helicity, i.e. on the small-scale velocity gradients.
In HIT, enstrophy is well known to be closely related to dissipation: they have the same mean value up to a constant given by the fluid viscosity $\nu$ (Tsinober, 2001). However, they are two different descriptors of the structure of the small scales of turbulence (Sreenivasan & Antonia, 1997), and extreme events of dissipation and enstrophy are rather very different. Extreme dissipation represents intense local strain, while large enstrophy represents strong vortical motions. Despite their intrinsic difference, however, Yeung et al. (2012) showed that at sufficiently large $Re$ extreme events of dissipation and enstrophy scale similar, and tend to occur together. Enstrophy shares some common features with energy as well, such as a transfer from large to small scales, with the dissipation being confined at the small scales. However, the conventional notion of a cascade does not apply for enstrophy: it is not inviscidly conserved because vortex stretching ($\bm{\omega}\cdot\bm{\nabla}\bm{u}$) acts as a source of vorticity at all scales (Davidson et al., 2008). Most of enstrophy is generated at the small scales and does not have to be transferred among space and scales to be dissipated. Nevertheless, vortex stretching is active also in the inertial range of scales (where dissipation is negligible), and this enstrophy has to be transferred from large to small scales to be eventually destroyed. The importance of vortex stretching in turbulence is well known since the pioneering works of Taylor (1938); Betchov (1956); Ashurst et al. (1987). Though, there is a lack of consensus of the role of vortex stretching in the energy cascade. Despite the earlier works (see for example Davidson, 2004; Davidson et al., 2008; Doan et al., 2018) providing evidence about the relation between energy cascade and vortex stretching, some recent works have shown that the energy cascade is mainly driven by the self-amplification of the strain-rate field, and that vortex stretching plays a key role only during fluctuations of the cascade about its average value (Carbone & Bragg, 2020).
Helicity, a pseudoscalar defined as $h=\bm{\omega}\cdot\bm{u}$, is another inviscid invariant of the three-dimensional ($3D$) Navier–Stokes equations. That is, like energy, helicity is also conserved by the non linearity of the Navier–Stokes equations (Moffatt, 1969). It admits topological interpretations in relation to the linkages of the flow vortex lines (Moffatt, 1969; Moffatt & Tsinober, 1992): it is related to the knottedness of the vorticity lines, and is a descriptor of the breaking of parity invariance (mirror symmetry). Unlike energy, however, it is not a sign-definite quantity. The invariance of helicity has been associated with the conservation of the linkages of the vortex lines that move with the flow (Moffatt, 1969), and it is based on two main notions, i.e. (i) vortex lines behave like material lines under evolution of the inviscid Navier–Stokes equations, and (ii) the flux of vorticity through any open surface bounded by a curve moving with the fluid is conserved. Based on the conservation of total helicity, Brissaud et al. (1973) were the first to envisage the possible simultaneous existence of energy and helicity cascades in $3D$ turbulence, similar to what happens in two-dimensional ($2D$) turbulence where the two invariants are the energy and the enstrophy (Boffetta & Ecke, 2012; Falkovich et al., 2017). Based on phenomenological arguments, they proposed that two different scenarios are possible, when dealing with helical flows. The first admits a simultaneous cascade of energy and helicity that leads to a $-5/3$ power-law spectrum for both quantities. The second scenario of a pure helicity cascade (i.e. no energy cascade) predicts a range of power-laws for both energy and helicity. Kraichnan (1973) argued that, unlike in $2D$ turbulence where the existence of invariant enstrophy effectively blocks the forward energy cascade, the possibility of a joint energy and helicity cascade is more plausible (see also André & Lesieur, 1977) in $3D$ turbulence. This was later confirmed by Polifke & Shtilman (1989), and by the numerical simulations of Borue & Orszag (1997). They observed a cascade of helicity from large to small scales, and did not detect an inverse cascade of helicity. In the following years, several authors studied helicity cascade in helical homogeneous isotropic turbulence. Chen et al. (2003b) found using direct numerical simulations that the energy and helicity fluxes feature a plateau in the inertial range of scales confirming the existence of a joint helicity and energy cascade from large to small scales. They also investigated intermittency in these fluxes and observed that the scaling exponents for the helicity fluxes are smaller compared to those for energy. The helicity flux is thus intrinsically more intermittent than the energy flux, consistent with the observation that helicity essentially behaves like a passive scalar (Sreenivasan & Antonia, 1997; Romano & Antonia, 2001). Over the last years, several authors have studied helicity cascades, mainly using the Fourier space statistics (see for example Biferale et al., 2013; Alexakis, 2017; Pouquet et al., 2019). Most of these works exploit an exact decomposition of the velocity field in a helical Fourier basis to properly account for triad interaction between wavenumbers (Constantin & Majda, 1988; Waleffe, 1992).
1.3 The Kármán-Howarth-Hill or the Generalised Kolmogorov Equation
It is therefore clear that to fully characterise the influence of the polymeric additives on the Kolmogorov picture of turbulence, one has to consider also enstrophy and helicity besides the turbulent kinetic energy. In this respect, Baj et al. (2022) introduced a generalised framework which is suitable for investigating the scale-space transfers of energy, enstrophy and helicity in Newtonian turbulent flows. Their framework extends the generalised Kolmogorov equation, or GKE, (Marati et al., 2004; Danaila et al., 2004; Cimarelli et al., 2013; Gatti et al., 2020; Gattere et al., 2023), also known as Kármán-Howarth-Monin-Hill or Kármán-Howarth-Hill equation (Alves Portela et al., 2017; Yasuda & Vassilicos, 2018; Alves Portela et al., 2020; Yao et al., 2024), introduced by Hill (2001, 2002). The GKE is a generalisation of the Kármán-Howarth equation (Frisch, 1995; Pope, 2000), and is directly derived from the Navier–Stokes equations without any assumptions; it does not require either isotropy nor homogeneity. The GKE is an exact budget equation for the second-order structure function, commonly referred to as scale energy (Davidson & Pearson, 2005), and characterises the mechanisms of production, transfer and dissipation of energy in the combined space of scales and positions. The GKE and its generalisations have been applied to several flow configurations over the years, mainly to study how inhomogeneity changes the Richardson and Kolmogorov picture of turbulence (Casciola et al., 2003; Cimarelli et al., 2016; Alves Portela et al., 2017; Mollicone et al., 2018; Cimarelli et al., 2021; Chiarini et al., 2022b, a; Apostolidis et al., 2023). De Angelis et al. (2005) used the GKE to investigate the influence of polymeric additives on the energy cascade in homogeneous isotropic turbulence, but at rather small Deborah numbers ($De\leq 0.5$), for which the influence of the polymeric stresses is subdominant at all scales.
1.4 The present study
In this work we investigate the influence of the polymeric additives on the scale transfers of energy, helicity and enstrophy in HIT at a relatively large Re and over a wide range of De. First, we extend the formulation of Baj et al. (2022) and provide the exact scale-by-scale budget equations for energy, helicity and enstrophy for polymer-laden turbulent flows. Compared to the purely Newtonian case, the new resulting budget equations feature additional terms that capture the influence of the polymers on the scale-space transfers, and detail the scale-space exchanges of energy, helicity and enstrophy between the fluid and polymeric phases. The new set of equations has been derived without any approximation, and are valid for a generic inhomogeneous and anisotropic turbulent flow. Then, we use these equations to elucidate the influence of the polymers on the production and transfers of energy, helicity and enstrophy in a homogeneous isotropic turbulent flow of dilute polymeric suspensions. New insights are provided, with a particular look at the influence of the polymers on the intermittent nature of the transfers, and on the local flow topology. The study relies on the database introduced by Singh & Rosti (2024) and obtained by means of direct numerical simulations. The Reynolds number is set at $Re_{\lambda}\approx 460$, while the Deborah number is varied in the $1/9\leq De\leq 9$ range.
The remainder of the work is structured as follows. In §2 we introduce the scale-space budget equations. In §2.3 the details of the DNS database used in this work are briefly recalled. Sections §4, §5 and §6 deal with our findings, and are respectively devoted to the scale-by-scale budget equations for energy, enstrophy and helicity. In §5 the effect of the polymers on the local flow topology is also discussed in relation with their influence on the vortex stretching. A concluding discussion and perspectives are then provided in §7.
2 Mathematical formulation and numerical method
2.1 The governing equations
The governing equations for an incompressible turbulent flow with polymeric additives are
$\begin{gathered}\frac{\partial u_{i}}{\partial t}+u_{j}\frac{\partial u_{i}}{\partial x_{j}}=-\frac{1}{\rho}\frac{\partial p}{\partial x_{i}}+\nu\frac{\partial^{2}u_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}}{\partial x_{j}}+f_{i},\\ \frac{\partial R_{ij}}{\partial t}+u_{k}\frac{\partial R_{ij}}{\partial x_{k}}=\frac{\partial u_{i}}{\partial x_{k}}R_{kj}+R_{ik}\frac{\partial u_{k}}{\partial x_{j}}-\frac{1}{\tau_{p}}\left(\mathcal{P}R_{ij}-\delta_{ij}\right),\\ \frac{\partial u_{j}}{\partial x_{j}}=0,\end{gathered}$ (1)
where $u_{i}$ is the velocity field, $p$ is the reduced pressure, $\rho$ is the fluid density, $\nu$ is the fluid kinematic viscosity, and $f_{i}$ is the external forcing used to sustain the flow. The presence of the polymers is accounted for in the momentum equation by means of the extra-stress tensor $T_{ij}$, which is related to the conformation tensor $R_{ij}$ as
$R_{ij}=T_{ij}\frac{\tau_{p}}{\mu_{p}}+\delta_{ij},$ (2)
where $\tau_{p}$ is the polymeric relaxation time, $\mu_{p}$ is the polymeric viscosity and $\delta_{ij}$ is the Kronecker delta. $\mathcal{P}$ is the Peterlin function and is equal to $\mathcal{P}=1$ for the purely elastic Oldroyd-B model and to $\mathcal{P}=(\ell_{max}^{2}-3)/(\ell_{max}^{2}-R_{ii})$ for the FENE-P model; $\ell_{max}$ is the maximum polymer extensibility, $R_{ii}$ is the instantaneous measure of the end-to-end length of the polymers.
By taking the curl of the momentum equation we obtain the equation for vorticity $\omega_{i}=\varepsilon_{ijk}\partial u_{k}/\partial x_{j}$ as,
$\frac{\partial\omega_{i}}{\partial t}+u_{j}\frac{\partial\omega_{i}}{\partial x_{j}}=\omega_{j}\frac{\partial u_{i}}{\partial x_{j}}-\nu\frac{\partial^{2}\omega_{i}}{\partial x_{j}\partial x_{j}}+\frac{1}{\rho}\frac{\partial T_{ij}^{\omega}}{\partial x_{j}}+f_{i}^{\omega},$ (3)
where $T_{i\ell}^{\omega}=\varepsilon_{ijk}\partial T_{k\ell}/\partial x_{j}$ , $f_{i}^{\omega}=\varepsilon_{ijk}\partial f_{k}/\partial x_{j}$, and $\varepsilon_{ijk}$ is the Levi-Civita symbol.
[…]
2.2 The budget equations
Following the work of Baj et al. (2022), we consider three specific structure functions, i.e. the velocity structure function $\delta q^{2}=\delta u_{i}\delta u_{i}$, the vorticity structure function $\delta\omega^{2}=\delta\omega_{i}\delta\omega_{i}$ and the helicity structure function $\delta h=\delta u_{i}\delta\omega_{i}$ where repeated indices imply summation. The three structure functions feature velocity and vorticity incremements ($\delta u_{i}$ and $\delta\omega_{i}$) between two points $\bm{x}$ and $\bm{x}^{\prime}$, that can be identified by means of their midpoint $\bm{X}=(\bm{x}+\bm{x}^{\prime})/2$ and separation vector $\bm{r}=\bm{x}^{\prime}-\bm{x}$, namely
$\delta u_{i}\left(\bm{X},\bm{r},t\right)=u_{i}\left(\bm{X}+\frac{\bm{r}}{2},t\right)-u_{i}\left(\bm{X}-\frac{\bm{r}}{2},t\right),$
and
$\delta\omega_{i}\left(\bm{X},\bm{r},t\right)=\omega_{i}\left(\bm{X}+\frac{\bm{r}}{2},t\right)-\omega_{i}\left(\bm{X}-\frac{\bm{r}}{2},t\right).$
In the most general case $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle$ depend upon seven independent variables $\bm{X}$, $\bm{r}$ and $t$.
The budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ describe production, transport and dissipation of energy, enstrophy and helicity in the space of scales $\bm{r}$ and positions $\bm{X}$. These equations link the variation in time of the three structure functions at a given scale and position to the instantaneous unbalance among production, transport and dissipation. The three budget equations are obtained after manipulation of the Navier–Stokes equations for the velocity and the vorticity, without any assumption of homogeneity and isotropy; for the full derivation (for a purely Newtonian fluid) of the equation for $\left\langle{\delta q^{2}}\right\rangle$ we refer the interested reader to the appendix of Gatti et al. (2020). The derivation of the budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ starts with the evolution equations for velocity and vorticity, and requires a sequence of manipulations that use the incompressibility constraint. A brief recap of the main steps is also provided in Baj et al. (2022).
In compact form, the evolution equations for $\left\langle{\delta q^{2}}\right\rangle=\left\langle{\delta u_{i}\delta u_{i}}\right\rangle$, $\left\langle{\delta h}\right\rangle=\left\langle{\delta u_{i}\delta\omega_{i}}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle=\left\langle{\delta\omega_{i}\delta\omega_{i}}\right\rangle$ can be written as:
$\frac{\partial\left\langle{\delta g}\right\rangle}{\partial t}+\frac{\partial\psi^{\delta g}_{j}}{\partial X_{j}}+\frac{\partial\phi^{\delta g}_{j}}{\partial r_{j}}=P^{\delta g}+\Pi^{\delta g}+\mathcal{E}^{\delta g}+F^{\delta g},$ (4)
where $\delta g$ is a generic quantity. On the left hand side, $\bm{\psi}^{\delta g}$ and $\bm{\phi}^{\delta g}$ are the components of the six-dimensional vector field of fluxes $\bm{\Phi}^{\delta g}$ in the physical space $\bm{X}$ and in the space of scales $\bm{r}$ respectively. On the right hand side, $P^{\delta g}$ is the production term, $\Pi^{\delta g}$ denotes the exchange of $\delta g$ between the fluid phase and the polymeric microstructure at position $\bm{X}$ and scale $\bm{r}$, $\mathcal{E}^{\delta g}$ is the fluid dissipation term, and $F^{\delta g}$ denote the energy entering the system at a certain position and scale due to the external forcing.
In the following, we detail the terms of the budget equation for the three structure functions. The components of the flux in the physical space $\bm{\psi}$ read
$\psi_{j}^{\delta q^{2}}=$ $\underbrace{\left\langle{U_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}+\underbrace{\frac{2}{\rho}\left\langle{\delta p\delta u_{j}}\right\rangle}_{\text{Pressure transp.}}$ (5)
$\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta T_{ij}\delta u_{i}}\right\rangle,}_{\text{Polymeric transp.}}$
$\psi_{j}^{\delta\omega^{2}}=\underbrace{\left\langle{U_{j}^{*}\delta\omega^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta\omega^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta T^{\omega}_{ij}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (6)
$\psi_{j}^{\delta h}=$ $\underbrace{\left\langle{U_{j}^{*}\delta h}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{u_{j}^{*}\delta h}\right\rangle-\frac{1}{2}\left\langle{\omega_{j}^{*}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}+\underbrace{\frac{1}{\rho}\left\langle{\delta p\delta u_{j}}\right\rangle}_{\text{Pressure transp.}}$ (7)
$\underbrace{-\frac{\nu}{2}\frac{\partial\left\langle{\delta h}\right\rangle}{\partial X_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{1}{2\rho}\left(\left\langle{\delta T^{\omega}_{ij}\delta u_{i}}\right\rangle+\left\langle{\delta T_{ij}\delta\omega_{i}}\right\rangle\right),}_{\text{Polymeric transp.}}$
where $j=1,2,3$ for the general case with three non homogeneous directions. The components of the flux vector in the space of scales $\bm{\phi}$ are:
$\phi_{j}^{\delta q^{2}}=\underbrace{\left\langle{\delta U_{j}\delta q^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{4}{\rho}\left\langle{T_{ij}^{*}\delta u_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (8)
$\phi_{j}^{\delta\omega^{2}}=\underbrace{\left\langle{\delta U_{j}\delta\omega^{2}}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta\omega^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{4}{\rho}\left\langle{T^{\omega*}_{ij}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (9)
$\phi_{j}^{\delta h}=\underbrace{\left\langle{\delta U_{j}\delta h}\right\rangle}_{\text{Mean transp.}}+\underbrace{\left\langle{\delta u_{j}\delta h}\right\rangle-\frac{1}{2}\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle}_{\text{Turbulent transp.}}\underbrace{-2\nu\frac{\partial\left\langle{\delta h}\right\rangle}{\partial r_{j}}}_{\text{Viscous diff.}}\underbrace{-\frac{2}{\rho}\left\langle{T_{ij}^{\omega*}\delta u_{i}}\right\rangle-\frac{2}{\rho}\left\langle{T_{ij}^{*}\delta\omega_{i}}\right\rangle,}_{\text{Polymeric transp.}}$ (10)
where $j=1,2,3$. The first terms on the right of equation 4 are the production terms,
$P^{\delta q^{2}}=\underbrace{-2\left\langle{u_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-2\left\langle{\delta u_{i}\delta u_{j}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*},}_{\text{Mean Prod. (Grad. $U$)}}$ (11)
$\begin{gathered}P^{\delta\omega^{2}}=\underbrace{-2\left\langle{u_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-2\left\langle{\delta u_{j}\delta\omega_{i}}\right\rangle\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad.$\Omega$)}}+\\ \underbrace{2\left\langle{\omega_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+2\left\langle{\delta\omega_{i}\delta\omega_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Production (Grad. $U$)}}+\\ \underbrace{2\left\langle{\delta\omega_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+2\left\langle{\delta\omega_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Mean Vortex Stretching}}+\\ \underbrace{2\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+2\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Turbulent Vortex Stretching}},\end{gathered}$ (12)
$P^{\delta h}=$ $\underbrace{-\left\langle{u_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)-\left\langle{\delta u_{j}\delta u_{i}}\right\rangle\left(\frac{\partial\Omega_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad. $\Omega$)}}+\underbrace{\left\langle{\delta u_{i}\Omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta u_{i}\delta\Omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle}_{\text{Mean Vortex stretching}}+$
$\underbrace{-\left\langle{u_{j}^{*}\delta\omega_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)-\left\langle{\delta u_{j}\delta\omega_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}+\left\langle{\omega_{j}^{*}\delta u_{i}}\right\rangle\delta\left(\frac{\partial U_{i}}{\partial x_{j}}\right)+\left\langle{\delta\omega_{j}\delta u_{i}}\right\rangle\left(\frac{\partial U_{i}}{\partial x_{j}}\right)^{*}}_{\text{Mean Prod. (Grad. $U$)}},$ (13)
and the fluid-polymer exchange term
$\begin{gathered}\Pi^{\delta q^{2}}=-\frac{4}{\rho}\underbrace{\left\langle{T_{ij}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\pi^{\delta q^{2}}}\right\rangle^{*}},\qquad\Pi^{\delta\omega^{2}}=-\frac{4}{\rho}\underbrace{\left\langle{T_{ij}^{\omega}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\pi^{\delta\omega^{2}}}\right\rangle^{*}}\qquad\text{and}\\ \Pi^{\delta h}=-\frac{4}{\rho}\underbrace{\left(\frac{1}{2}\left\langle{T_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}+\frac{1}{2}\left\langle{T_{ij}^{\omega}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}\right).}_{\left\langle{\pi^{\delta h}}\right\rangle^{*}}\end{gathered}$ (14)
The last terms on the right hand side correspond to the dissipation
$\mathcal{E}^{\delta q^{2}}=-4\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{*}},\ \mathcal{E}^{\delta\omega^{2}}=-4\underbrace{\nu\left\langle{\frac{\partial\omega_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle^{*}},\ \text{and}\ \mathcal{E}^{\delta h}=-4\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle^{*}}_{\left\langle{\varepsilon_{f}^{\delta h}}\right\rangle^{*}},$ (15)
and to the external forcing
$F^{\delta q^{2}}=2\left\langle{\delta f_{i}\delta u_{i}}\right\rangle,\ F^{\delta\omega^{2}}=2\left\langle{\delta f^{\omega}_{i}\delta\omega_{i}}\right\rangle,\ \text{and}\ F^{\delta h}=\left\langle{\delta f^{\omega}_{i}\delta u_{i}}\right\rangle+\left\langle{\delta f_{i}\delta\omega_{i}}\right\rangle.$ (16)
In the previous equations, the asterisk superscript $\cdot^{*}$ denotes the average between the positions $\bm{X}\pm\bm{r}/2$, i.e. $a^{*}=(a(\bm{X}+\bm{r}/2)+a(\bm{X}-\bm{r}/2))/2$.
The various terms appearing in the expressions of the fluxes are interpreted as the mean and the turbulent transport, pressure transport and viscous diffusion. Note that pressure does not play a role for $\bm{\psi}^{\delta\omega^{2}}$ due to the curl operator. As already stated in Baj et al. (2022), the budget equation for $\left\langle{\delta h}\right\rangle$ shows that the transfer of helicity in the $\bm{X}$ and $\bm{r}$ spaces associated with turbulence fluctuations arise not only from the interaction of $\delta h$ with the velocity increments $\delta u_{j}$, but also from the interaction of $\delta q^{2}$ with the vorticity increments $\delta\omega_{j}$. Accordingly, there result two different routes of helicity transfer, similar to the results of Yan et al. (2020). The presence of polymers introduces an additional contribution to $\bm{\psi}$ and $\bm{\phi}$, identified as polymeric transport. The right hand side of equation 4 describes the net source of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ in space and among scales. The mean production terms identify how $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ are exchanged with the mean flow. Accordingly, similar to the single-point budget equations, the production terms are related with the mean-flow gradients. Unlike for $\left\langle{\delta q^{2}}\right\rangle$, the mean production of $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle$ depends on both $\partial U_{i}/\partial x_{j}$ and $\partial\Omega_{i}/\partial x_{j}$. Likewise, enstrophy production receives additional contributions due to vortex stretching by turbulent fluctuations. When present, the mean vorticity gradient also contributes to vortex stretching, and therefore to the production of $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta\omega^{2}}\right\rangle$. In addition to the fluid dissipation terms $\mathcal{E}$, the right hand side of equation 4 features an additional sink/source $\Pi$ which describes the exchange of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ with the polymeric microstructure. This term is not positive-definite, meaning that exchanges in either direction are possible locally.
2.2.1 Homogeneous isotropic turbulence
For homogeneous flows, the three structure functions and the terms of the corresponding budget equations lose their dependence on $\bm{X}$, and are functions of only $\bm{r}\equiv(r_{x},r_{y},r_{z})$. Exploiting isotropy, we integrate in the $\bm{r}$ space over spherical shells of radius $r=|\bm{r}|$, with surface $S(r)$ and volume $\Omega(r)$, and obtain the generalised Kármán-Howarth equation for homogeneous isotropic polymeric turbulence. The three relations respectively for the velocity ($\left\langle{\delta q^{2}}\right\rangle$), vorticity ($\left\langle{\delta\omega^{2}}\right\rangle$) and helicity ($\left\langle{\delta h}\right\rangle$) structure functions are given as follows, i.e.
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta q^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f}^{\delta q^{2}}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{4}{\rho}\left\langle{\delta u_{i}T_{ij}^{*}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta q^{2}}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta q^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta q^{2}}}\right\rangle(r)}\\ -\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle}r-\frac{4}{3}\underbrace{\frac{1}{\rho}\left\langle{T_{ij}\frac{\partial u_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\pi^{\delta q^{2}}}\right\rangle}r+\underbrace{\frac{1}{S(r)}\int_{V(r)}2\left\langle{\delta u_{i}\delta f_{i}}\right\rangle\text{d}V}_{\left\langle{F^{\delta q^{2}}}\right\rangle(r)},\end{gathered}$ (17)
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta\omega^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f}^{\delta\omega^{2}}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{4}{\rho}\left\langle{\delta\omega_{i}T^{\omega*}_{ij}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta\omega^{2}}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta\omega^{2}}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta\omega^{2}}}\right\rangle(r)}\\ +\underbrace{\frac{1}{S(r)}\int_{V(r)}2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right)\text{d}V}_{\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle(r)}-\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial\omega_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle}r\\ -\frac{4}{3}\underbrace{\frac{1}{\rho}\left\langle{T^{\omega}_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\pi^{\delta\omega^{2}}}\right\rangle}r+\underbrace{\frac{1}{S}(r)\int_{V(r)}2\left\langle{\delta f^{\omega}_{i}\delta\omega_{i}}\right\rangle\text{d}V}_{\left\langle{F^{\delta\omega^{2}}}\right\rangle(r)}\end{gathered}$ (18)
and
$\begin{gathered}\underbrace{\frac{1}{S(r)}\oint_{S(r)}\left\langle{\delta u_{j}\delta h}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f,a}^{\delta h}}\right\rangle(r)}+\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{1}{2}\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle n_{j}\text{d}S}_{\left\langle{\Phi_{f,b}^{\delta h}}\right\rangle(r)}\\ +\underbrace{\frac{1}{S(r)}\oint_{S(r)}-\frac{2}{\rho}\left(\left\langle{T_{ij}^{\omega*}\delta u_{i}}\right\rangle+\left\langle{T_{ij}^{*}\delta\omega_{i}}\right\rangle\right)n_{j}\text{d}S}_{\left\langle{\Phi_{p}^{\delta h}}\right\rangle(r)}=\underbrace{\frac{1}{S(r)}\oint_{S(r)}2\nu\frac{\partial\left\langle{\delta h}\right\rangle}{\partial r_{j}}n_{j}\text{d}S}_{\left\langle{D_{f}^{\delta h}}\right\rangle(r)}\\ -\frac{4}{3}\underbrace{\nu\left\langle{\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial\omega_{i}}{\partial x_{j}}}\right\rangle}_{\left\langle{\varepsilon_{f}^{\delta h}}\right\rangle}r-\frac{4}{3}\underbrace{\left\langle{\frac{1}{2\rho}\left(T_{ij}^{\omega}\frac{\partial u_{i}}{\partial x_{j}}+T_{ij}\frac{\partial\omega_{i}}{\partial x_{j}}\right)}\right\rangle}_{\left\langle{\pi^{\delta h}}\right\rangle}r+\underbrace{\frac{1}{S(r)}\int_{V(r)}\left(\left\langle{\delta f^{\omega}_{i}\delta u_{i}}\right\rangle+\left\langle{\delta f_{i}\delta\omega_{i}}\right\rangle\right)\text{d}V}_{\left\langle{F^{\delta h}}\right\rangle(r)}.\end{gathered}$ (19)
These equations reveal the main differences between Newtonian and polymeric turbulence: for Newtonian fluids, energy is dissipated by only the fluid viscosity $\varepsilon_{f}^{\delta q^{2}}$, and the flux of energy is only due to the nonlinear term $\Phi_{f}^{\delta q^{2}}$, alongside the assumption that in the inertial range of scales $D_{f}^{\delta q^{2}}$ and $F^{\delta q^{2}}$ are negligible. In this range, a (positive) production term associated with vortex stretching arises only for the enstrophy equation. For polymeric flows, the total flux comprises an additional component $\Phi_{p}$ that mimics the energy transfer among scales associated with the polymeric microstructure. An additional source/sink term $\pi$ appears on the right hand side, that characterises the exchange of energy, helicity and enstrophy between the fluid and the polymers. It is interesting to note that for the energy, $\langle\pi^{\delta q^{2}}\rangle$ matches the polymeric dissipation $\langle\epsilon_{p}\rangle\equiv\left\langle{\mu_{p}(R_{ii}-3)/(2\tau_{p}^{2})}\right\rangle$, meaning that on average this term is a net sink for the fluid phase (De Angelis et al., 2005).
When studying the scale-by-scale energy budget of polymeric turbulence in Fourier space, Rosti et al. (2023) showed that the non Newtonian flux is not a purely dissipative term. They isolate the purely dissipative part of this flux by assuming that it has the same asymptotic dependence as the fluid viscous flux at small scales. An additional requirement is that at vanishing scales it matches the polymeric dissipation $\langle\pi^{\delta q^{2}}\rangle=\langle\varepsilon_{p}\rangle$. Here we use the same approach for the three budgets. We thus split the polymeric flux $\Phi_{p}$ into two contributions,
$\Phi_{p}=\Phi_{p,i}+D_{p},\ \text{with}\ \left\langle{D_{p}}\right\rangle=\frac{\left\langle{\pi}\right\rangle}{\left\langle{\varepsilon_{f}}\right\rangle}\left\langle{D_{f}}\right\rangle\ \text{and}\ \left\langle{\Phi_{p,i}}\right\rangle=\left\langle{\Phi_{p}}\right\rangle-\left\langle{D_{p}}\right\rangle.$
To simplify the notation, hereafter we use $\Phi_{p}$ instead of $\Phi_{p,i}$ to indicate the inertial contribution to the polymeric flux.
2.3 The numerical database
This work relies on the DNS database introduced by Singh & Rosti (2024). Their paper contains full details on the numerical method and the related computational procedures, which are only briefly recalled here. Equations 1 are numerically integrated using the in-house solver Fujin (https://www.oist.jp/research/research-units/cffu/fujin), which uses an incremental pressure-correction scheme. The Navier–Stokes equations written in primitive variables are solved on a staggered grid using second-order finite-differences in all the directions. The momentum equation is advanced in time using a second-order Adams-Bashforth time scheme, while the non Newtonian stress tensor is advanced in time with a second-order Crank-Nicolson scheme. A log-conformation formulation (Izbassarov et al., 2018) ensures positive-definiteness of the conformation tensor at all times. Turbulence is sustained using the Arnold-Beltrami-Childress (ABC) cellular-flow forcing (Podvigina & Pouquet, 1994).
The equations are solved within a cubic domain of size $L=2\pi$ having periodic boundary conditions in all directions, discretised with $N^{3}=1024^{3}$ grid points, to ensure that all the scales down to the smallest dissipative ones are properly resolved, i.e. $\eta/\Delta x=O(1)$, where $\Delta x$ is the grid spacing and $\eta$ the Kolmogorov scale. The parameters are chosen to achieve a Taylor-microscale Reynolds number of $Re_{\lambda}=u^{\prime}\lambda/\nu\approx 460$ in the purely Newtonian case ($u^{\prime}$ is the root mean square of the velocity fluctuations and $\lambda$ is the Taylor length scale). The Deborah number $De=\tau_{p}/\tau_{f}$, where $\tau_{f}=L/u_{rms}$ is the large-eddy turnover time, is varied in the range $1/9\leq De\leq 9$. For all cases, the fluid and polymer viscosities are fixed such that $\mu_{f}/(\mu_{f}+\mu_{p})=0.9$. Simulations are advanced with a constant time step of $\Delta t/\tau_{\eta}=2\times 10^{-3}$, where $\tau_{\eta}$ is the Kolmogorov time scale. Details of the numerical simulations with bulk quantities of interest are provided in table 1.
Table 1: Details of the numerical simulations considered in the present study. $De$ is the Deborah number, $Re_{\lambda}$ is the Reynolds number based on $u^{\prime}=\sqrt{2E/3}$ and on the Taylor length scale $\lambda$, $\eta$ is the Kolmogorov length scale, $R_{ii}$ is the trace of the conformation tensor and mimics the free energy of the polymeric phase, $\varepsilon_{f}^{\delta q^{2}}$, $\varepsilon_{f}^{\delta\omega^{2}}$ and $\varepsilon_{f}^{\delta h}$ are the fluid dissipation of energy, enstrophy and helicity, and $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$, $\pi^{\delta h}$ are the source/sink of energy, enstrophy and helicity due to the fluid-polymer coupling.
$De$ $Re_{\lambda}$ $\left\langle{\lambda}\right\rangle$ $\left\langle{\eta}\right\rangle$ $\langle\varepsilon_{f}^{\delta q^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle$ $\langle\varepsilon_{f}^{\delta h^{2}}\rangle$ $\langle\pi^{\delta q^{2}}\rangle$ $\langle\pi^{\delta\omega^{2}}\rangle$ $\langle\pi^{\delta h}\rangle$
-- $458.6$ $0.1578$ $0.0037$ $59.82$ $5.1041\times 10^{5}$ $61.51$- -- --
$1/9$ $513.64$ $0.1747$ $0.0039$ $50.84$ $3.6950\times 10^{5}$ $55.14$ $6.50$ $2.5279\times 10^{4}$ $5.51$
$1/3$ $627.06$ $0.2037$ $0.0041$ $41.76$ $2.4098\times 10^{5}$ $44.38$ $17.29$ $2.0524\times 10^{4}$ $12.76$
$1$ $795.80$ $0.2767$ $0.0050$ $21.04$ $1.1265\times 10^{5}$ $33.09$ $39.94$ $-7.0836\times 10^{4}$ $33.84$
$3$ $761.11$ $0.2859$ $0.0053$ $15.50$ $9.1522\times 10^{4}$ $25.21$ $33.52$ $-7.7998\times 10^{4}$ $22.48$
$9$ $776.30$ $0.2733$ $0.0050$ $22.03$ $1.2596\times 10^{5}$ $33.06$ $45.61$ $-1.0975\times 10^{5}$ $36.42$
3 Structure functions and dissipation
Figure 1: Dependence of the velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ (left), vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ (centre) and helicity structure function $\left\langle{\delta h}\right\rangle$ (right) on the Deborah number. For $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ the lines for the different $De$ are vertically shifted for increase the clarity.
We start probing the velocity, vorticity and helicity structure functions $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ (see figure 1), and discuss how they are influenced by polymeric additives.
The velocity structure function $\left\langle{\delta q^{2}}\right\rangle$ clearly shows the multiscaling behaviour of polymeric flows (Zhang et al., 2021; Rosti et al., 2023; Singh & Rosti, 2024). In the Newtonian case, and for $De\leq 1/3$, we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$ in the intermediate range of scales in agreement with the Kolmogorov predictions (Kolmogorov, 1941). For small $De$, i.e. in the limit of $\tau_{p}\rightarrow 0$, the polymers do not stretch, and they only marginally influence the velocity structure function. For $De\approx 1$ the polymeric relaxation time is comparable with the large time scale of the flow and the polymers interact with the energy cascade, inducing a steeper depletion of $\left\langle{\delta q^{2}}\right\rangle$ with $r$ in the so-called elastic range of scales (see for example De Angelis et al., 2005; Perlekar et al., 2010). For $De=1$ we observe $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ that matches the energy spectrum scaling of $\kappa^{-2.3}$ (Rosti et al., 2023), where $\kappa$ is the wavenumber. Due to large relaxation times, the polymers now remain stretched when further increasing $De$ and begin to decouple from the carrier flow. For $De\geq 3$, therefore, the influence of the polymers weakens, and $\left\langle{\delta q^{2}}\right\rangle$ shows a less steeper dependence on $r$, progressively recovering the $r^{2/3}$ Kolmogorov scaling. For $De=9$ we measure $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.1}$ over a relatively wide range of scales. Note that, $\left\langle{\delta q^{2}}\right\rangle\sim r^{2}$ at the small scales $r\rightarrow 0$ for all cases, which is known to follow from the smoothness of the velocity field at small scales (Schumacher et al., 2007). With the real space structure function, indeed, we can not observe the increase of the energy in the deep dissipative range that is found in the Fourier space energy spectrum by Perlekar et al. (2010), Rosti et al. (2023), and by Singh & Rosti (2024) with the present database.
The vorticity structure function $\left\langle{\delta\omega^{2}}\right\rangle$ is shown in the central panel of figure 1. It is clear that enstrophy growth (generation) is confined to small scales for all $De$, as already observed by other authors for HIT (see for example Jiménez et al., 1993; Ishihara et al., 2013; Elsinga et al., 2017). Indeed, $\left\langle{\delta\omega^{2}}\right\rangle$ rapidly increases with $r$ at the small scales and saturates in the inertial range. As detailed in the following, the polymers reduce the vorticity fluctuations at the small scales (Liberzon et al., 2005, 2006; Perlekar et al., 2010) which results in a decrease of $\left\langle{\delta\omega^{2}}\right\rangle$ at all $r$. Note that, like for $\left\langle{\delta q^{2}}\right\rangle$, the effect of the polymers on the vorticity structure function is non-monotonous, with $\left\langle{\delta\omega^{2}}\right\rangle$ being minimum for $De=3$ at all scales.
The dependence of the helicity structure functions $\left\langle{\delta h}\right\rangle$ on $De$ is shown in the right panel of figure 1. In HIT, we find $\left\langle{\delta h}\right\rangle\sim r^{2/3}$ in the inertial range, in agreement with the existence of the cascade of helicity from larger to the smaller scales envisaged by Brissaud et al. (1973); see also Polifke & Shtilman (1989), Borue & Orszag (1997) and Baj et al. (2022) and section §6. Note also that $\left\langle{\delta h}\right\rangle$ and $\left\langle{\delta q^{2}}\right\rangle$ have an almost identical $r^{2/3}$ range, indicating that the inertial range for helicity has an almost identical span across scales as that for energy, confirming the findings of Chen et al. (2003a). The effect of the polymers on $\left\langle{\delta h}\right\rangle$ resembles what has been observed for $\left\langle{\delta q^{2}}\right\rangle$, suggesting that the two cascades are influenced in a similar fashion. For small and large $De$, the influence of the polymers is rather small: the $r^{2/3}$ scaling is recovered for $De\rightarrow 0$ and $De\rightarrow\infty$. For intermediate $De$, instead, $\left\langle{\delta h}\right\rangle$ shows a steeper dependence on $r$ in the elastic range of scales. For $De=1$, we find $\left\langle{\delta h}\right\rangle\sim r^{0.9}$, which is shallower compared to the $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ scaling for the energy, hinting a weaker influence of polymers on the helicity cascade. Overall, the results show that helicity is mainly concentrated at the large scales for all $De$: turbulence cascade tends to restore the mirror symmetry at the small scales in both Newtonian and polymeric turbulence (Ditlevsen & Giuliani, 2001; Chen et al., 2003a; Baj et al., 2022); see §6 for additional discussion.
To provide further insights on the scale-by-scale influence of the polymers on the energy and enstrophy distributions, we consider the signature function $V(r)$, introduced by Davidson (2004) and Davidson & Pearson (2005) to eliminate the enstrophy contribution of eddies of scales larger than $r$ from the second-order structure function, i.e.
$V(r)=-\frac{1}{2}r^{2}\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial}{\partial r}\left(\frac{3}{4}S_{2}(r)\right)\right).$
The signature function has the following properties: (i) $\int_{0}^{r}V(r)\text{d}r\geq 0$ and (ii) $\int_{0}^{\infty}V(r)\text{d}r=\frac{1}{2}\left\langle{\bm{u}^{2}}\right\rangle$(Davidson, 2004). Therefore, $V(r)$ is interpreted as the energy density and $rV(r)$ estimates the kinetic energy associated with eddies of scale $r$. Similarly, $10V(r)/r$ is a measure of the enstrophy associated with eddies of size $r$.
Figure 2: Contribution of eddies of size $r$ on the (left) total energy and on the (centre) total enstrophy for different Deborah numbers. $V(r)$ is the signature function introduced by Davidson (2004) and Davidson & Pearson (2005). The right panel shows the distribution of the helicity $h=\bm{u}\cdot\bm{\omega}$ for different $De$.
The left and middle panels of figure 2 show the scale-by-scale distributions of $rV(r)$ and $10V(r)/r$ for different $De$. We start with the energy distribution in the left panel. As expected, largest scales are the most energetic, and $rV(r)$ diminishes as $r$ decreases, being null for $r\rightarrow 0$. The influence of the polymers is felt more at the small scales, while the amount of energy associated with the largest scales remains largely unaffected. This is consistent with the observations in earlier studies (Bhattacharjee & Thirumalai, 1991; Fouxon & Lebedev, 2003; De Angelis et al., 2005; Perlekar et al., 2010) that polymers decrease the energy content at small and intermediate scales, with $r_{p}$, that separates the large and intermediate scales, changing with $De$. In agreement with the findings of Rosti et al. (2023), the energy depletion at the small scales does not show a monotonous dependence on $De$. Interestingly, the effect of $De$ on $rV(r)$ changes with $r$, suggesting that the local coupling between the fluid and the polymers depends on the local ratio between the polymeric relaxation time and the characteristic time scale of the flow at that scale, i.e. $\tau_{p}/\tau_{f}(r)$, where $\tau_{f}(r)$ decreases at smaller scales (for example, K41 implies that the flow time scales in Newtonian turbulence decay as $\tau_{f}(r)\sim r^{2/3}$). To understand this better, let us focus on $1\leq De\leq 9$. At $r\approx 0.02$ the energy depletion is maximum for $De=3$ and minimum for $De=9$, while for $r\approx 0.6$ it is maximum for $De=9$ and minimum for $De=1$: the inversion of the trend occurs first at the smaller scales, where indeed the local $\tau_{p}/\tau_{f}(r)$ is larger and the decoupling between the fluid and the polymers occurs for smaller values of $\tau_{p}$. A last comment regards the overshoot of $rV(r)$, i.e. the gentle local increase of energy observed for scales between the inertial and the far-dissipation ranges, usually referred to as bottleneck (see for example Falkovich, 1994; Lohse & Müller-Groeling, 1995; Frisch et al., 2013). Although a detailed investigation of this effect is not within the scope of the present work, it is worth noticing that when the coupling between the fluid and the polymers is maximum, i.e. $De\approx 1$, this overshoot is annihilated and the local increase of $rV(r)$ is not observed.
We now briefly move to the distribution of $10V(r)/r$, shown in the central panel of figure 2. It confirms that the enstrophy is limited to the small scales for all $De$, and that the influence of the polymers in negligible at the largest scales. Akin to energy, and in agreement with the distribution of $\left\langle{\delta\omega^{2}}\right\rangle$, we observe that the influence of the polymers on the scale-by-scale enstrophy is not monotonic in $De$: the maximum enstrophy reduction is observed for $De=3$ at the small scales and for $De=9$ at larger scales.
The right panel of figure 2 describes the influence of the polymers on the probability distribution function (pdf) of $h=\bm{u}\cdot\bm{\omega}$. Helicity is not positive definite, and has an asymmetric distribution for all cases; it is right skewed and the right tail is longer for all $De$. In agreement with the distribution of $\left\langle{\delta h}\right\rangle$ shown in figure 1, the presence of the polymers narrows the distribution of $h$, suggesting that polymers reduce the extreme events of helicity. However, the asymmetry of the distribution increases, and even more so as $De$ increases, indicating that the presence of the polymers favours events that break mirror symmetry. Again, the $De$ dependence is not monotonous, with an inversion of the effects of the polymers on the tails observed for $De\geq 3$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Distribution of the dissipation of (left) energy, (centre) enstrophy, and (right) helicity for different Deborah numbers. The bottom panels plot the distribution of the three quantities normalised with their root-mean-square value.
We show the nature of dissipation of all three quantities in figure 3. These quantities play a key role in the scale-by-scale cascades. In Newtonian turbulence, for the two averaged, inviscid invariants (energy and helicity) the dissipation rate matches the injection rate at the large scales and the averaged total flux in the inertial range of scales. The distributions of the dissipations show large tails, which are indicative of small-scale intermittency (Sreenivasan & Antonia, 1997). As expected, the distributions narrow for the polymeric case: polymers reduce the magnitude of the velocity derivatives, thus largely reducing the regions of large dissipations (Liberzon et al., 2005, 2006). The distributions of the energy and enstrophy dissipation collapse nicely for all $De$ once the quantities are normalised with the respective standard deviation, similarly to what was shown by Perlekar et al. (2010) and is consistent with the $De$ invariance of intermittency corrections observed in Rosti et al. (2023).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 4: Distribution of the the polymer/fluid exchange terms $\pi$ for $\delta q^{2}$, $\delta\omega^{2}$ and $\delta h$ for different Deborah numbers. Left: $\pi^{\delta q^{2}}$, centre: $\pi^{\delta\omega^{2}}$ and right: $\pi^{\delta h}$. The bottom panels plot the distribution of the three quantities normalised with their root-mean-square value.
We conclude this section with the distributions of $\pi^{\delta q^{2}}$, $\pi^{\delta\omega^{2}}$ and $\pi^{\delta h}$, which capture the additional sink/source of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ due to the polymeric microstructure. These quantities are not positive definite, meaning that locally polymers may act as either a sink or a source for the velocity/vorticity fluctuations. The distribution of $\pi^{\delta q^{2}}$ is right-skewed for all $De$, with the highly right skewed distribution confirming the positive average value $\langle\pi^{\delta q^{2}}\rangle=\langle\epsilon_{p}\rangle$ (De Angelis et al., 2005). As $De$ increases, both the right and left tails widen, meaning that sink and source intense events are equally promoted by the larger stretching of the polymers. The effect of $De$ is much more significant on $\pi^{\delta\omega^{2}}$, whose distributions are shown in the central panels of figure 4. They become increasingly left skewed with increasing $De$. At small $De<1$, the distributions are right skewed and, on average, the polymers act as a sink for the enstrophy. For $De\geq 1$, instead, the distribution of $\pi^{\delta\omega^{2}}$ is left skewed. In this case, due to the larger stretching of the polymers, the fluid-polymer interaction facilitates intense production of enstrophy. As detailed in §5, this qualitative change is rooted in a profound re-organisation of the small-scale fluctuations. Eventually, the right panels in figure 4 shows that the distribution of $\pi^{\delta h}$ is symmetric, and that both tails become longer as $De$ increases.
4 Scale energy
In this section, we detail the influence of the polymeric additives on the scale-by-scale energy transfer. We first focus on the budget equation for $\left\langle{\delta q^{2}}\right\rangle$ and look at the modifications of the average energy cascade. This is the $r$-space equivalent of the Fourier analysis by Rosti et al. (2023). Then, we look in detail at the nonlinear ($\Phi_{f}^{\delta q^{2}}$) and polymeric ($\Phi_{p}^{\delta q^{2}}$) fluxes to highlight how polymers modify the intermittent nature of the energy transfers.
4.1 Scale-by-scale budget
Figure 5: Scale-by-scale energy budget. The panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$, from top-left to bottom-right. The yellow and blue shaded regions identify the inertial and elastic ranges of scales.
Figure 5 plots the scale-by-scale energy budget for both HIT and PHIT. We start looking at the HIT case ($De=0$), which depicts the Richardson-Kolmogorov picture of turbulence (Richardson, 1922). Energy is injected at the largest scales $r>r_{e}\approx 1.84$ by the external forcing $\text{d}\langle F^{\delta q^{2}}\rangle/\text{d}r\approx(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle$, and is then transferred from larger to smaller scales within the inertial range $0.043\approx r_{d}\leq r\leq r_{e}$ via the classical cascade process; $r_{e}$ denotes the scale above which the forcing term dominates (i.e. separates the energetic and inertial ranges of scales), while $r_{d}$ denotes the scale below which the dissipative term dominates (i.e. separates the inertial and dissipative ranges of scales). As predicted by the Kármán-Howarth equation (de Kármán & Howarth, 1938), the energy transfer rate is an invariant in the inertial range of scales and equals the dissipation rate $\text{d}\langle\Phi_{f}^{\delta q^{2}}\rangle/\text{d}r\approx-(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle$; see the pleateau in the top left panel in figure 5. At the smallest scales $r<r_{d}$ the viscous contribution $D_{v}^{\delta q^{2}}$ dominates, and $\left\langle{\delta q^{2}}\right\rangle$ is dissipated into heat by viscous friction.
Polymeric additives introduce an alternative route for energy transfer among scales ($\Phi_{p}^{\delta q^{2}}$) and an additional dissipative process ($D_{p}^{\delta q^{2}}$). For small $De$($<1$), the global picture resembles what observed in HIT. In this case, indeed, the polymer chains are not effectively stretched as their relaxation time is small compared to the characteristic time scales of the flow, and they quickly relax back to their equilibrium lengths; here, $\Phi_{p}^{\delta q^{2}}\approx D_{p}^{\delta q^{2}}\approx\pi^{\delta q^{2}}\approx 0$. As $De$ increases ($De\geq 1$) the relevance of the non-Newtonian contribution increases, and the polymeric additives modify the energy cascade in a non trivial way. Energy enters the system at the large scales ($r>r_{e}$) and is dissipated away at the same rate, via both the fluid and polymeric dissipation, i.e.
$\frac{\text{d}}{\text{d}r}\left\langle{F^{\delta q^{2}}}\right\rangle\approx\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}+\pi^{\delta q^{2}}}\right\rangle.$
In the intermediate range of scales ($r_{d}<r<r_{e}$) energy is then transferred from larger to smaller scales by two alternative routes, respectively associated with (i) the classic inertia-driven energy cascade $\Phi_{f}^{\delta q^{2}}$ and (ii) the fluid/polymer interaction $\Phi_{p}^{\delta q^{2}}$. Here the viscous effects are negligible and the total energy transfer rate is an invariant and matches the total dissipation rate, i.e.
$\frac{\text{d}}{\text{d}r}\left(\left\langle{\Phi_{p}^{\delta q^{2}}}\right\rangle+\left\langle{\Phi_{f}^{\delta q^{2}}}\right\rangle\right)\sim-\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}+\pi^{\delta q^{2}}}\right\rangle.$
Eventually, at the smallest scales ($r<r_{d}$) energy is dissipated away by both the fluid and the polymers, and
$\frac{\text{d}}{\text{d}r}\left\langle{D_{f}^{\delta q^{2}}}\right\rangle\rightarrow\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle\ \text{and}\ \frac{\text{d}}{\text{d}r}\left\langle{D_{p}^{\delta q^{2}}}\right\rangle\rightarrow\frac{4}{3}\left\langle{\pi^{\delta q^{2}}}\right\rangle\ \text{for}\ r\rightarrow 0.$
The range of scales where the energy cascade is effectively modified by the presence of the polymers changes with $De$; see the blue shaded region in figure 5. One can identify for each $De$ the scale $r_{p}^{*}$ below which elastic effects are expected to dominate using simple arguments. We estimate this scale by comparing the local turnover time of eddies of size $r$, i.e. $\tau_{f}(r)\sim r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ — where $\delta V(r)=(u_{i}(\bm{x}+\bm{r})-u_{i}(\bm{x}))\cdot r_{i}/r)$ — with the polymeric relaxation time $\tau_{p}$. Thus, $r_{p}^{*}$ is estimated as:
$\frac{r_{p}^{*}}{\sqrt{\left\langle{\delta V^{2}}\right\rangle^{*}}}=\tau_{p},\quad\text{where}\quad\left\langle{\delta V^{2}}\right\rangle^{*}=\left\langle{\delta V^{2}}\right\rangle(r_{p}^{*}).$ (20)
For $r>r_{p}^{*}$ the time scale of the fluid fluctuations is larger than $\tau_{p}$: the polymers are only marginally stretched and thus their influence on the flow is weak. On the contrary, when $r<r_{p}^{*}$ the time scale of the fluid is smaller than the polymeric relaxation time: the polymers are stretched more effectively at these scales and thus modify the organisation of the corresponding fluctuations. For the present cases, we measure $r_{p}^{*}\approx 0.04,0.32,1.34$ and $2.81$ for $De=1/9,1/3,1$ and $3$: the range of scales where elastic effects significantly modify the energy cascade increases with $De$, in agreement with the widening of the elastic range of scales shown in figure 5. For $De=9$, $\tau_{p}>r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ for all $r$, suggesting that the polymeric additives effectively modify the energy cascade in the whole range of scales. We reiterate that when $De$ is very large and $\tau_{p}\gg r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ the polymeric chains decouple from the carrier phase and their influence on the carrier flow is rather low.
The multiple scaling behaviour observed in figure 1 for $De\geq 1$ is consistent with $r_{p}^{*}$ being well within the $r_{e}<r<r_{d}$ range of scales (where both the viscous effects and the external forcing are negligible) for large enough Deborah and Reynolds numbers. The energy cascade driven by the fluid inertia $\Phi_{f}^{\delta q^{2}}$ dominates at the larger scales $r_{p}^{*}\lessapprox r\lessapprox r_{e}$, where indeed $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$. The cascade driven by the fluid/polymer interaction $\Phi_{p}^{\delta q^{2}}$, instead, dominates at the smaller $r_{d}\lessapprox r\lessapprox r_{p}^{*}$ scales, where $\left\langle{\delta q^{2}}\right\rangle\sim r^{\xi}$ with $\xi\approx 1.3\approx 4/3$. Accordingly, this peculiar multiscaling behaviour is not observed for $De<1$ and/or small $Re$, as in these cases $r_{p}^{*}$ is within the dissipative range of scales (see for example De Angelis et al., 2005; Perlekar et al., 2010).
A last comment regards the influence of $De$ on the dissipative range of scales. Our data show that $r_{d}$ (i.e. scale below which the total dissipative term $D_{v}^{\delta q^{2}}+D_{p}^{\delta q^{2}}$ dominates) decreases with $De$, thus indicating that the dissipative range of scales progressively shrinks as the polymeric relaxation time increases.
4.2 The nonlinear and non Newtonian fluxes
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 6: Dependence of the distribution of the nonlinear flux $\Phi_{f}$ on the Deborah number, for (left) $r=0.1$, (centre) $r=0.3$, and (right) $r=0.9$. Top: distribution of $\Phi_{f}^{\delta q^{2}}$. Bottom: distribution of $\Phi_{f}^{\delta q^{2}}/\Phi_{f,rms}^{\delta q^{2}}$. In the bottom panels the inset provide a zoom of the normalised distributions.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Dependence of the distribution of the non Newtonian flux $\Phi_{p}$ on the Deborah number, for (left) $r=0.1$, (centre) $r=0.3$, and (right) $r=0.9$. Top: distribution of $\Phi_{p}^{\delta q^{2}}$. Bottom: distribution of $\Phi_{p}^{\delta q^{2}}/\Phi_{p,rms}^{\delta q^{2}}$. In the bottom panels the inset provide a zoom of the normalised distributions.
The average picture of the cascade and the interscale exchanges in the $r-$space is however not representative of the actual physical processes. In the previous section, we have shown the influence of the polymers on the energy cascade in an average sense only, even though it is known that the interscale energy transfer is highly intermittent (see for example Piomelli et al., 1991; Domaradzki et al., 1993; Cerutti & Meneveau, 1998). Moreover, even though energy cascade in HIT is on an average from larger to smaller scales, there exist localised regions in the flow where energy actually cascades from smaller to larger scales, opposite to the average sense. We now investigate how the presence of polymeric additives modifies this picture, and whether this localised inverse energy transfer is also detected in the transfer mechanism driven by the fluid/polymer interaction.
Figures 6 and 7 plot the probability distribution functions of $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$ for different $De$. We consider three different separations $r$, i.e. $r=0.1$ where $\Phi_{p}^{\delta q^{2}}$ dominates for $De>1$, and $r=0.3,0.9$ where $\Phi_{f}^{\delta q^{2}}$ dominates. In the present convention, a positive $\Phi^{\delta q^{2}}>0$ indicates backscatter, i.e. local events where energy goes from smaller to larger scales, while a negative $\Phi^{\delta q^{2}}<0$ indicates events where energy goes from larger to smaller scales. The distributions reveal the highly intermittent and non-Gaussian nature of both $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$, given their heavy tails (see Ishihara et al., 2009; Yasuda & Vassilicos, 2018, for HIT): backscatter and extreme forward events exist with much higher probability than for a normal distribution. The asymmetry of the distributions shows that on average both $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$ are negative, and that the average interscale transfer of energy is in the forward sense for both routes. To be quantitative, for $De=1$ and $r=0.9$ ($r=0.1$) the probability of $\Phi_{f}^{\delta q^{2}}>\Phi^{\delta q^{2}}_{f,rms}$ is $0.0672$ ($0.0274$), while the probability of events with $\Phi_{f}^{\delta q^{2}}<-\Phi_{f,rms}^{\delta q^{2}}$ is $0.1640$ ($0.0767$); similarly, the probability of $\Phi_{p}^{\delta q^{2}}>\Phi_{p,rms}^{\delta q^{2}}$ is $0.0323$ ($0.0105$), while the probability of events with $\Phi_{p}^{\delta q^{2}}<-\Phi_{p,rms}^{\delta q^{2}}$ is $0.2402$ ($0.1545$).
We now focus on how polymers influence the tails of the distributions, i.e. the extreme events. We start looking at the nonlinear flux $\Phi_{f}^{\delta q^{2}}$ (see figure 6). For the three chosen scales, the addition of polymers leads to narrower tails: extreme events of both direct and inverse energy transfers along the classical route are inhibited. This effect becomes more evident as $De$ increases, in agreement with a smoother velocity field, and with the larger amount of energy that is redirected towards the polymer-driven transfer route. When looking at the fluxes normalised with their root-mean-square value (bottom panels), we note that the distributions at $r=0.3$ and $r=0.9$ collapse reasonably well for all $De$. This means that in the $\Phi_{f}^{\delta q^{2}}$ dominated range, the presence of the polymers modifies the strength of the extreme cascade events, but does not influence its intermittency. In other words, in the range of scales where the polymeric contribution is weak the self-similarity of the classical cascade process is preserved. At the small scales $r=0.1$ where $\Phi_{p}^{\delta q^{2}}$ dominates, however, the normalised distributions of $\Phi_{f}^{\delta q^{2}}$ do not overlap: here polymers modify the intermittency of the inertia-driven cascade. The nonlinear flux $\Phi_{f}^{\delta q^{2}}$ has a small average contribution at this scale, whose intermittency increases with $De$. For very large elasticity, the tails begin to shrink again as polymers decouple from these small scales easily thus reducing their effect on small-scale fluctuations.
Figure 7 considers the distribution of the polymeric flux $\Phi_{p}^{\delta q^{2}}$ at different scales. An increase of $De$ results in wider tails, with the asymmetry of the distribution increasing with $De$ in agreement with an increase of the average value of $\Phi_{p}^{\delta q^{2}}$. When considering the normalised flux (bottom panels), the scenario changes compared to $\Phi_{f}^{\delta q^{2}}$. In this case, indeed, the distributions overlap reasonably well at the small scale $r=0.1$, but not for the larger $r=0.3$ and $r=0.9$ scales. An increase of $De$ modifies the space-time intermittency of the interscale energy transfer driven by the polymeric microstructure at large scales only, while self-similarity with $De$ is observed at small $r$ in the $\Phi_{p}^{\delta q^{2}}$ dominated range. Accordingly with the increased asymmetry, the larger $De$ corresponds to a narrower positive tails and to a slightly heavier negative tail.
Figure 8: Distribution of $\cos(\theta)=\bm{u}^{*}\cdot(\delta\bm{\omega}\times\delta\bm{u})/(||\bm{u}^{*}||||\delta\bm{u}\times\delta\bm{\omega}||)$ for different scales. In order from top-left to bottom-right, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
We now provide a further insight on the influence of the polymers on the inertia-driven energy cascade process by looking at the simple relation introduced by Baj et al. (2022) that links the nonlinear flux with the Lamb vector $\delta\bm{\omega}\times\delta\bm{u}$. Exploiting the Lamb decomposition and manipulating the equations, it is indeed possible to show that
$\frac{\partial}{\partial r_{j}}\left\langle{\delta u_{j}\delta q^{2}}\right\rangle=2\left\langle{\bm{u}^{*}\cdot(\delta\bm{\omega}\times\delta\bm{u})}\right\rangle.$ (21)
Since $||\delta\bm{u}\times\delta\bm{\omega}||^{2}=||\delta\bm{u}||^{2}||\delta\bm{\omega}||^{2}-||\delta h||^{2}$, it is clear that there is an implicit connection between the scale-by-scale helicity $\delta h$ and the interscale energy transfer. Indeed, a large $\delta h=\delta\bm{\omega}\cdot\delta\bm{u}$ corresponds to a small $\delta\bm{\omega}\times\delta\bm{u}$, and therefore to a weaker local transfer of energy. This agrees with the observation of previous authors which found that the magnitude of the helicity has an impact on the local transfer of energy (see for example Pelz et al., 1985; Stepanov et al., 2015). To gain further insights on how polymers influence the interscale energy transfer, we look at the angle $\theta$ between $\bm{u}^{*}$ and $\delta\bm{\omega}\times\delta\bm{u}$ such that
$\cos(\theta)=\frac{\bm{u}^{*}\cdot\left(\delta\bm{\omega}\times\delta\bm{u}\right)}{||\bm{u}^{*}||||\delta\bm{u}\times\delta\bm{\omega}||}.$ (22)
Figure 8 plots the distributions of $\cos(\theta)$ at different scales $r$ for $0\leq De\leq 9$. We start our discussion with HIT, where the distribution is positively skewed for all $r$. This is consistent with equation 21, which implies that on average the left hand side ($||\bm{u}^{*}||||\delta\bm{\omega}\times\delta\bm{u}||\cos(\theta)$) is positive. Similar to what found by Baj et al. (2022), as the separation increases the probability of parallel ($\cos(\theta)=1$) and antiparallel ($\cos(\theta)=-1$) alignment increases, with a larger probability of the parallel events. For PHIT ($De>0$), the picture changes in a scale-dependent manner. At small $De(\ll 1)$, as expected, the distributions remain similar to HIT. For large $De(\gtrsim 1)$, instead, at small scales $r$ where the polymeric flux $\Phi_{p}^{\delta q^{2}}$ dominates the Lamb vector has a larger likelihood of being normal to the local advecting velocity, with the distribution still being right skewed. This agrees with the fact that in this range of scales, only a small fraction of energy is transferred to smaller scales by the inertia-driven transfer route. The weakening of the $\text{d}\Phi_{f}^{\delta q^{2}}/\text{d}r$ interscale-energy transfer in the elastic range of scales is thus accompanied by an increase of the events where $\bm{u}^{*}(r)$ and $\delta\bm{\omega}(r)\times\delta\bm{u}(r)$ are perpendicular. When considering large $r$, where the influence of the polymers is weaker and $\Phi_{f}^{\delta q^{2}}$ dominates, the distribution of $\cos(\theta)$ flattens and progressively recovers a distribution similar to that in HIT.
5 Scale enstrophy
5.1 Scale-by-scale budget
Figure 9: Scale-by-scale enstrophy budget. In order from left to right and from top to bottom, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
We plot the scale-by-scale budget for $\left\langle{\delta\omega^{2}}\right\rangle$ in figure 9. Unlike energy and helicity, enstrophy is not an inviscid invariant of the Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, that quantifies the amount of enstrophy that is created up to the scale $r$ by means of vortex stretching-like processes. This means that even in HIT, the notion of cascade used for energy transfer does not apply here, and most of the enstrophy is generated directly at the small scales. However, as stated by several authors (see for example Davidson et al., 2008), vortex stretching does occur in the inertial range of HIT and hence a net transfer of enstrophy to smaller scales is expected at these $r$.
We start looking at HIT ($De=0$). Figure 9 shows that in the inertial range, vortex stretching is balanced by the dissipation of enstrophy, i.e., $\langle V_{st}\rangle(r)\approx(4/3)\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, resulting in the $\langle V_{st}\rangle(r)\sim r^{1}$ scaling. This can be easily seen by rewriting equation 18 for Newtonian turbulence, by dropping the polymeric contributions. In this range of scales the forcing $\langle F^{\delta\omega^{2}}\rangle$ and dissipation $\langle D_{f}^{\delta\omega^{2}}\rangle$ contributions to the total flux $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ disappear, and $\langle\Phi_{f}^{\delta\omega^{2}}\rangle\ll\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle$ as shown in figure 9: we indeed have that
$\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle(r)\approx-\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle r.$ (23)
A more insightful way to understand such a linear scaling behaviour of $\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle$ is by realising that vortex stretching receives its maximum contribution from the smallest scales. This can be seen by the dimensional estimate of the scale-by-scale vorticity $\omega_{r}$
$\omega_{r}\sim\frac{u_{r}}{r}\sim\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}r^{-2/3},$ (24)
which becomes maximum as $r\to\eta$, so that $\omega_{\eta}\sim\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}\eta^{-2/3}$. Now, to estimate the cumulative vortex stretching $\left\langle{V_{st}}\right\rangle$ up to scale $r$,
$\left\langle{V_{st}}\right\rangle(r)$ $=\frac{1}{S(r)}\int_{V(r)}\text{d}V\ 2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right),$ (25)
we begin by estimating the scale-by-scale vortex stretching $v^{st}_{r}$. This is given by the terms within the integral sign in the above equation 25, i.e.,
$v^{st}_{r}\sim\omega_{r}\omega_{r}\frac{u_{\eta}}{\eta}.$ (26)
Indeed, $v^{st}_{r}$ also gets a maximal contribution from the smallest scales (see also Davidson et al., 2008) which is given as
$v^{st}_{\eta}\sim\omega^{2}_{\eta}\ \frac{u_{\eta}}{\eta}\sim\left(\frac{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}}{\eta^{2/3}}\right)^{3}\sim\frac{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle}{\eta^{2}}\sim\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle,$ (27)
where we have used the $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle/\eta^{2}$ dimensional estimate. Ignoring the contribution to vortex stretching from larger scales, we can estimate the cumulative vortex stretching upto a scale $r$ as $\langle V_{st}\rangle(r)\approx 1/S(r)\int_{V(r)}\text{d}Vv^{st}_{\eta}\sim r^{-2}r^{3}\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, which is consistent with the data plotted in figure 9. Moreover, the nonlinear flux shows a $\langle\Phi_{f}^{\delta\omega^{2}}\rangle(r)\sim r^{-1}$ scaling and its magnitude increases as $r$ decreases within the inertial range. Since enstrophy is mainly generated at the small scales, its transfer rate among scales strengthens as $r$ decreases, with a maximum close to the transition from the inertial to the dissipative range of scales. We follow Davidson et al. (2008) to use dimensional analysis to explain the observed scaling law for $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$. The idea is that in the inertial range of scales the enstrophy cascade is based on mechanisms that are similar to vortex stretching. For example, let us consider a vortex tube of length $\ell$ and cross-rection $A\sim r^{2}$. Under the action of the external mean shear, the vortex tube is stretched and its cross-section decreases. This means enstrophy at a larger scale has been converted to a more intense enstrophy at a smaller scale. Based on this assumption, the transfer of enstrophy among scales in the inertial range is driven by eddies with scale $r$ and velocity $\delta u\sim u_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{1/3}$. When dealing with $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$, we are concerned with the scale-by-scale vorticity $\omega_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{-2/3}$ to obtain the observed scaling law, i.e.,
$\left\langle{\Phi_{f}^{\delta\omega^{2}}}\right\rangle(r)\sim\frac{1}{S(r)}\int_{S(r)}\text{d}S\ u_{r}\omega^{2}_{r}n_{j}\sim r^{-2}r^{2}\ \left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{1/3}r^{-4/3}=\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{-1}\sim r^{-1}.$ (28)
Refer to captionRefer to captionRefer to caption
Figure 10: (left) Dependence of the Vortex stretching term on $r$ for different $De$. (right) Dependence of $\left\langle{\pi^{\delta\omega^{2}}}\right\rangle/\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle$ on $De$.
We now discuss how the presence of polymers influence the scale-by-scale budget for $\left\langle{\delta\omega^{2}}\right\rangle$. The addition of polymers gives three additional terms in the budget equation, i.e., $\langle\pi^{\delta\omega^{2}}\rangle$ that describes the exchange of enstrophy between the polymeric and fluid phases, and $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle D_{p}^{\delta\omega^{2}}\rangle$ that represent the inertial and dissipative enstrophy flux contribution resulting from the fluid/polymer interaction. In other words, like for $\left\langle{\delta q^{2}}\right\rangle$, an additional source/sink and alternative mechanisms for the enstrophy scale transfer and dissipation arise. For small $De<1$, both $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle\pi^{\delta\omega^{2}}\rangle$ are rather small, and the classic enstrophy production $\langle V_{st}\rangle$ and transfers $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ are only marginally influenced by the polymers; see the top right and centre left panels of figure 9. For these $De$, $\langle\pi^{\delta\omega^{2}}\rangle<0$, meaning that on average the polymers act as a sink of enstrophy for the fluid phase; see also figure 10. For large Deborah numbers the scenario is different. As $De$ increases, the intensity of $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ only slightly changes, indicating that the polymers only marginally affect the classical transfer of enstrophy from larger to smaller scales. However, for $De>1$ the intensity of $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle D_{p}^{\delta\omega^{2}}\rangle$ equals that of $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ and $\langle D_{f}^{\delta\omega^{2}}\rangle$: for large elasticity, the two fluid- and polymer-driven scale-by-scale transfer and dissipative processes have comparable relevance at all scales $r$. This differs from what is seen for $\left\langle{\delta q^{2}}\right\rangle$, where the global energy cascade is dominated by the nonlinear flux at the larger scales and by the polymeric flux at the smaller ones.
The vortex stretching is strongly modulated for $De>1$. Figure 10 shows that $\langle V_{st}\rangle(r)$ largely weakens as $De$ increases, accordingly with the experimental results of Liberzon et al. (2005) and Liberzon et al. (2006). To be quantitative, we measure $\langle V_{st}\rangle/\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle/r|_{r=L/2}\approx 1.1$ for $De=0$ and $\langle V_{st}\rangle/\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle/r|_{r=L/2}\approx 0.35$ for $De=9$. Notably, the weakening of the vortex stretching is accompanied by a change of sign of $\langle\pi^{\delta\omega^{2}}\rangle$, that for $De>1$ becomes positive and balances the viscous dissipation; see figures 4 and 10. In other words, at large $De$ the polymers annihilate the classical production of enstrophy via vortex stretching, but its interaction with the fluid phase results into a new net production of enstrophy at all scales. As discussed in the following section, this change of behaviour at large $De$ can be explained with a change of the local flow topology.
5.2 Local strain
In this subsection, we relate the influence of the polymeric additives to the vortex stretching $V_{st}(r)$ by looking at the local topology of the flow. The local topology of the flow can be described using the three principle invariants of the velocity gradient tensor $A_{ij}\equiv\partial u_{j}/\partial x_{i}$ (Davidson, 2004; Meneveau, 2011). $A_{ij}$ can be decomposed into its symmetric and antisymmetric parts, i.e. the rate-of-strain tensor $S_{ij}$ and the rate-of-rotation tensor $W_{ij}$
$A_{ij}=\frac{\partial u_{i}}{\partial x_{j}}=\underbrace{\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)}_{S_{ij}}+\underbrace{\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{j}}{\partial x_{i}}\right)}_{W_{ij}}.$ (29)
$A_{ij}$ depends on the three principal rates of strain $\alpha\geq\beta\geq\gamma$ and the three components of the vorticity $\bm{\omega}$ along the principal axes. In other words, the local topology of the flow is entirely determined by (i) the three principal rates of strain $\alpha,\beta$ and $\gamma$, (ii) the magnitude of the vorticity, i.e. the enstrophy $\omega^{2}=\bm{\omega}\cdot\bm{\omega}$, and (iii) the orientation of $\bm{\omega}$ relative to the principal axes of strain. Additionally, the incompressibility condition implies that the sum of the eigenvalues of $S_{ij}$, $\alpha+\beta+\gamma=0$. This also means that at least one eigenvalue is not positive ($\gamma\leqslant 0$), while at least one must be non-negative ($\alpha\geqslant 0$). The statistics of vortex stretching are related to those of the eigenvalues of $S_{ij}$ as (see Davidson, 2004)
$\left\langle{\omega_{i}\omega_{j}S_{ij}}\right\rangle=-4\left\langle{\alpha\beta\gamma}\right\rangle.$ (30)
In HIT, $04\left\langle{\alpha\beta\gamma}\right\rangle=-\left\langle{\omega_{i}\omega_{j}S_{ij}}\right\rangle<$ 0 implying $\beta>0$. That is, on an average, we have one large compressive strain is accompanied by two weaker extensional ones. It is known that in HIT, on an average $(\left\langle{\alpha}\right\rangle,\left\langle{\beta}\right\rangle,\left\langle{\gamma}\right\rangle)\approx(3,1,-4)|\beta|$ (Davidson, 2004; Meneveau, 2011). Note that, although this seems to be consistent with the generation of sheet-like structures, it is also consistent with the stretching of vortex tubes, once the self-induced strain is considered (Davidson, 2004).
We now look at the influence of the polymers on the distribution of $\alpha$, $\beta$ and $\gamma$ to relate the $V_{st}(r)$ with the different extension/compression of the fluid elements.
Figure 11: Eigenvalues of $S_{ij}$. (top) Distribution of $s^{*}=-3\sqrt{6}\alpha\beta\gamma/(\alpha^{2}+\beta^{2}+\gamma^{2})^{3/2}$ for different $De$. (bottom) Distribution of (left) $\alpha$, (centre) $\beta$, and (right) $\gamma$ for different $De$.
To this end, we consider the ratio
$s^{*}=-\frac{3\sqrt{6}\alpha\beta\gamma}{(\alpha^{2}+\beta^{2}+\gamma^{2})^{3/2}},$ (31)
that was introduced by Lund & Rogers (1994). For a random velocity gradient field with no preferred structure, the distribution of $s^{*}$ is uniform. Note that $s^{*}=1$ is equivalent to $\alpha=\beta=-\gamma/2>0$, and corresponds to a state of axisymmetric extension such that a small spherical fluid element moving in the flow extends (symmetrically) in two directions and contracts in the third. An instance of $s^{*}=0$ corresponds to $\beta=0$, meaning that the state of the straining is two-dimensional and the fluid element is equally stretched/compressed in two directions. In this case, $\alpha\beta\gamma=0$ and the production of enstrophy is null $\omega_{i}\omega_{j}S_{ij}=0$.
The top panel of figure 11 plots the distribution of $s^{*}$ for $0\leq De\leq 9$. As shown by Lund & Rogers (1994), for HIT the distribution peaks at $s^{*}=1$, confirming that the most likely state of straining is an axisymmetric extension. In PHIT, instead, the distribution of $s^{*}$ progressively flattens as $De$ increases, and the most probable value moves towards $s^{*}=0$, i.e. towards a two-dimensional state. The larger preference for $s^{*}=0$ (quasi-two-dimensionality) and not $s^{*}=-1$ is confirmed by the separate distributions of $\alpha$, $\beta$ and $\gamma$ in the bottom panels of figure 11. For the Newtonian case, the results confirm what is found with $s^{*}$, with $\beta$ being mostly positive and $(\left\langle{\alpha}\right\rangle,\left\langle{\beta}\right\rangle,\left\langle{\gamma}\right\rangle)=(4.3,1,-5.3)|\left\langle{\beta}\right\rangle|$. For PHIT, the distribution of the three eigenvalues shrinks and their most likely value monotonically decreases with $De$. To be quantitative, we quote the average values in table 2. This means that in HIT axisymmetric extensions are favoured, while in PHIT shear and planar extensional regions are more frequent. This agrees with the weakening of $V_{st}(r)$ and is consistent with the results by Warwaruk & Ghaemi (2024) that investigated the local flow topology of a polymer-laden turbulent boundary layer. Moreover, these results also validate the suggested mechanism of polymer drag reduction in polymer-laden wall-bounded flows provided for example by Lumley (1973) and Roy et al. (2006). They observed that the extensional viscosity of the polymers strongly inhibits uniaxial and biaxial flow regions, thus favouring two-dimensional states of strain and mitigating the formation of quasi-streamwise vortices.
$De$ $\left\langle{\alpha}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\gamma}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\beta}\right\rangle$
$1/9$ $4.3$ $-5.3$ $11$
$1/3$ $4.5$ $-5.5$ $12$
$1$ $5.9$ $-6.9$ $5$
$3$ $7.5$ $-8.5$ $4$
$9$ $9.5$ $-10.5$ $3.5$
Table 2: Dependence of the average eigenvalues of $S_{ij}$ on $De$.
Figure 12: (left) $QR$ and (right) $Q_{S}Q_{W}$ maps for (blue) $De=0$ and (red) $De=9$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 13: Distribution of (left) $Q$ and (right) $R$ for different $De$.
To determine the topology of the flow motion, we focus on two of the invariants of $A_{ij}$, i.e. $Q$ and $R$ (Cantwell, 1993), and plot their distributions in figures 12 and 13. A second-order tensor in three-dimensions possesses three invariants, which are directly related to its eigenvalues $\lambda_{i}$ by means of the characteristic equation
$\lambda^{3}-P\lambda^{2}+Q\lambda+R=0.$ (32)
The three invariants are
$P=\lambda_{1}+\lambda_{2}+\lambda_{3}=tr(A)=\alpha+\beta+\gamma\ (=0\ \text{for incompressible flow}),$ (33)
$Q=tr(A^{2})=(\lambda_{1}\lambda_{2}+\lambda_{2}\lambda_{3}+\lambda_{3}\lambda_{1})=-(\alpha^{2}+\beta^{2}+\gamma^{2})+\frac{\omega^{2}}{4},$ (34)
and
$R=\det(A)=-(\lambda_{1}\lambda_{2}\lambda_{3})=-\alpha\beta\gamma-\frac{\omega_{i}\omega_{j}S_{ij}}{4},$ (35)
where repeated indices are summed over and $tr(A)$ denotes the trace of the generic tensor $A$. The second invariant $Q$ measures the relative strength of strain and vorticity, with $Q\ll 0$ indicating regions of strong strain, and $Q\gg 0$ indicating regions of intense vorticity. The third invariant is a measure of the relative intensity of the production of vorticity ($R<0$) and strain ($R>0$). In figure 12 we also plot the discriminant $\Delta=27R^{2}/4+Q^{3}=0$ ($P=0$) curve for equation 32 with a black line. Below this curve, $\lambda_{i}$s are all real and the flow is dominated by strain. Above this curve, however, $A_{ij}$ has one real and two complex eigenvalues, and enstrophy dominates the flow. When $Q$ is large and positive, strain is locally weak and $R\sim-\omega_{i}\omega_{j}S_{ij}$: in this case $R<0$ implies vortex stretching, while $R>0$ implies vortex compression. When instead $Q$ is large and negative then $R\sim-\alpha\beta\gamma$: a negative $R$ implies a region of axial strain ($\alpha>0$; $\beta,\gamma<0$), while a positive $R$ implies a region of biaxial strain ($\alpha,\beta>0$; $\gamma<0$).
In HIT, the $Q-R$ joint distribution has a tear-drop shape with a clear point at the right-Vieillefosse tail with $\Delta=0$, $R>0$ and $Q<0$ (Ooi et al., 1999; Elsinga & Marusic, 2010). The largest probability is of events lying in the two quadrants where $QR<0$, meaning that there is strong negative correlation between $Q$ and $R$. In other words, the two most common states are vortex stretching $\omega_{i}\omega_{j}S_{ij}>0$ and biaxial strain $\alpha\beta\gamma<0$ (Betchov, 1956; Davidson, 2004). The distributions of $Q$ and $R$ become narrower in PHIT, in agreement with similar observations at small $Re$ by Perlekar et al. (2010). This means that the presence of polymers inhibits the occurrence of vortical and dissipative motions as well as of intense fluid extensions and compressions. This is conveniently visualised in figures 12 and 13 by the shrinking distributions of $R$ with increasing $De$. The $Q-R$ joint distribution for $De=9$ shows that there is still a bias for biaxial extensions in PHIT ($\alpha,\beta,-\gamma,-Q,R>0$). However, stretching is largely diminished compared to HIT, in agreement with more frequent two-dimensional states. The left panel of figure 13 shows that an increase in $De$ leads to a more symmetric distribution of $Q$ showing a stronger inhibition of events with positive $Q$ (large enstrophy). Moreover, the shrinking of the $Q>0$ tail indicates weaker vorticity while the shrinking of the tails of the $R$ distribution hint at a weaker vortical stretching. This is consistent with the discussion in preceding sections.
Figure 14: Qualitative influence of the polymers on the local structure of turbulence. Isocontours of $Q=3Q_{\text{rms}}$ for (top left) $De=0$, (top right) $De=1/3$, (bottom left) $De=1$, and (bottom right) $De=9$ from instantaneous snapshots.
We corroborate these observations with a qualitative picture of the influence of the polymers on the local structure of the flow: see figure 14 where the isocontours of $Q$ are plotted in instantaneous snapshots. For the $De=1$ case, where the fluid-polymer coupling is maximum, the flow features thicker and more elongated vortical structures with more vacant $Q\leq 3Q_{\text{rms}}$ regions. This agrees with the narrower $Q$ distribution discussed above.
The right panel of figure 12 shows the $Q_{S}-Q_{W}$ joint distribution, where $Q_{S}$ and $Q_{W}$ are the second invariants of the $S_{ij}$ and $W_{ij}$ tensors such that $Q_{S}+Q_{W}=Q$:
$Q_{S}=-\frac{1}{2}tr(S^{2})\ \text{and}\ Q_{W}=-\frac{1}{2}tr(W^{2}).$ (36)
$Q_{S}$ and $Q_{W}$ respectively capture the local rates of strain and rotation. These invariants are related to fluid dissipation $\varepsilon_{f}^{\delta q^{2}}$ and fluid enstrophy $\omega^{2}$ as
$\varepsilon_{f}^{\delta q^{2}}=-4\nu Q_{S}\ \text{and}\ \omega^{2}=4Q_{W}.$
Therefore, the $Q_{S}-Q_{W}$ joint distribution also indicates whether the flow is dominated by dissipation or enstrophy. We look at $\mathcal{K}=(-Q_{W}/Q_{S})^{1/2}$ (Truesdell, 1954); when $\mathcal{K}=0$ the flow is extension dominated ($Q_{S}\gg Q_{W})$, when $\mathcal{K}=\infty$ the flow undergoes rigid rotation locally, and is vorticity dominated ($Q_{S}\ll Q_{W}$), when $\mathcal{K}=1$ rotation and stretching are equal, as typical for vortex sheets and shear layers. In HIT, events with $Q_{W}>-Q_{S}$ are more frequent, meaning that the flow is mainly dominated by rigid rotations. In PHIT, instead, the shape of the distribution changes and events with $Q_{W}=-Q_{S}$ ($\mathcal{K}=1$) are favoured. This is in line with the above observation that polymers favour two-dimensional strain states. Note that the influence of polymers on the distribution of $Q_{S}$ ($-\varepsilon_{f}^{\delta q^{2}}$) and $Q_{W}$ ($\omega^{2}$) is rather different. This is visualised in the bottom left panel of figure 13, which considers $Q/Q_{rms}$, with $Q_{S}$ and $Q_{W}$ contributing to the negative and positive values of $Q$, respectively. When $De$ increases, indeed, the positive tail of $Q/Q_{rms}$ ($Q_{W}$) narrows, while the negative one ($Q_{S}$) becomes wider: the presence of polymers suppresses extreme rigid rotation events more than extreme extensions.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 15: Distribution of the cosine of the angle between the vorticity and the eigenvectors of the strain-rate tensor, for different $De$.
As said above, the orientation of $\bm{\omega}$ with respect to the principal axes of $S_{ij}$ is needed to completely determine the local flow topology. We therefore look at the influence of the polymers on the local orientation between the vorticity $\omega$ and $S_{ij}$ (Ashurst et al., 1987; Meneveau, 2011). Figure 15 plots the alignment between the vorticity $\bm{\omega}$ and the principal axes of the strain rate tensor $S_{ij}$. Our results show that polymers favour the anti-alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\alpha}$ (note that the $\bm{\omega}\cdot\hat{\bm{e}}_{\alpha}=0$ peak monotonically increases with $De$), and favour the alignemnt between $\bm{\omega}$ and $\hat{\bm{e}}_{\beta}$ (note that the $\bm{\omega}\cdot\hat{\bm{e}}_{\beta}=\pm 1$ peak monotonically increases with $De$).
6 Scale helicity
6.1 Scale-by-scale budget
Figure 16: Scale-by-scale helicity budget. In order from top-left to bottom-right, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
We now move to the scale-by-scale helicity budget. In the present set up, turbulence is sustained with the ABC forcing that injects both energy and helicity at the largest scales. As such, like energy, helicity is expected to cascade from large to small scales in the inertial range of scales, and to eventually vanish at the small scales where the mirror symmetry $\bm{u}\cdot\bm{\omega}=0$ is restored.
Our results give evidence of a dual direct cascade of energy and helicity for all $De$ (see figure 16), in agreement with the predictions by Kraichnan (1973), Polifke & Shtilman (1989) and Borue & Orszag (1997) for HIT. For the purely Newtonian case, $\text{d}\langle F^{\delta h}\rangle/\text{d}r\sim(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ at large scales $r>r_{e}$, $\text{d}\langle\Phi_{f}^{\delta h}\rangle/\text{d}r\sim(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ in the inertial range of scales $r_{d}<r<r_{e}$, and $\text{d}\langle D_{f}^{\delta h}\rangle/\text{d}r\rightarrow(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ for $r\rightarrow 0$. Similar to what was observed for $\left\langle{\delta q^{2}}\right\rangle$, polymers provide an additional transfer and dissipative mechanism for helicity, which become more and more important as $De$ increases. Analogous to the cascade of $\left\langle{\delta q^{2}}\right\rangle$, at the intermediate scales the helicity cascade is mainly driven by the nonlinear flux $\langle\Phi_{f}^{\delta h}\rangle$ at scales $r_{p}^{*}<r<r_{e}$ and by the polymeric flux $\langle\Phi_{p}^{\delta h}\rangle$ at scales $r_{d}<r<r_{p}^{*}$. The helicity cascade driven by inertia $\langle\Phi_{f}^{\delta h}\rangle$ is indeed progressively depleted as $De$ increases and the polymers properly stretch, similarly to what was observed for $\langle\Phi_{f}^{\delta q^{2}}\rangle$. Note that the existence of two different cascade processes that take over in two distinct range of scales is consistent with the multiscaling behaviour of $\left\langle{\delta h}\right\rangle$ reported in figure 1.
Refer to captionRefer to captionRefer to caption
Figure 17: Dependence of $\langle\Phi_{f,a}^{\delta h}\rangle$ (left) and $\langle\Phi_{f,b}^{\delta h}\rangle$ (right) on $r$ for different $De$.
Figure 17 provides a closer look at the influence of the polymeric additives on $\langle\Phi_{f}^{\delta h}\rangle$. As mentioned in §2, the nonlinear flux associated with the helicity consists of two distinct contributions. Besides the usual flux term $\langle\Phi_{f,a}^{\delta h}\rangle$ that gives a transfer due to velocity fluctuations $\left\langle{\delta h\delta u_{j}}\right\rangle$, $\langle\Phi_{f}^{\delta h}\rangle$ features an additional term $\langle\Phi_{f,b}^{\delta h}\rangle$ that describes a helicity flux due to $\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle$. Figure 17 shows that $\langle\Phi_{f,a}^{\delta h}\rangle$ dominates in HIT, being almost twice of $\langle\Phi_{f,b}^{\delta h}\rangle$ at all scales, with helicity being mainly transferred among scales by means of the classic convective term. When increasing $De$, instead, both the contributions to the flux progressively weaken in the elastic range of scales, meaning that polymers deplete both processes while redirecting part of the helicity transfer via the new polymer-driven flux. Notably, polymers have a stronger influence on $\langle\Phi_{f,a}^{\delta h}\rangle$, such that for $De=9$ the two processes on an average contribute equally to the transfer ($\langle\Phi_{f,a}^{\delta h}\rangle\approx\langle\Phi_{f,b}^{\delta h}\rangle$).
6.2 Relative helicity
Refer to captionRefer to captionRefer to caption
Figure 18: (left) Dependence of the relative helicity $r\left\langle{\delta h}\right\rangle/(2\left\langle{\delta q^{2}}\right\rangle)$ on $r$ for different $De$. (right) Dependence of $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle$ on $r$ for different $De$.
In this subsection, we look at the relative helicity defined as $h_{r}\equiv r\left\langle{\delta h}\right\rangle/(2\left\langle{\delta q^{2}}\right\rangle)$ after Borue & Orszag (1997), to assess how the polymers modify the tendency of the flow to restore mirror symmetry at the smallest scales. Figure 18 shows that the flow is maximally helical at the large scales, and that the relative helicity decays as $r$ decreases. Helicity is indeed dynamically unimportant at small scales. For HIT, we find that $h_{r}\sim r^{1}$ in the inertial range, as also shown by Borue & Orszag (1997). This is visualised in the right panel of figure 18, where $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle\sim 0.75$, which is similar to the corresponding quantity measured for a passive scalar (see Borue & Orszag, 1997). For PHIT, we find that $h_{r}$ decays at a smaller rate at intermediate $r$ compared to HIT, and that the $r^{1}$ scaling is lost in the inertial range as $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ scale there differently for $De>0$ as already shown in figure 1. We find that $h_{r}$ increases with $De$ at all scales. Overall, in agreement with the distribution of $h=\bm{u}\cdot\bm{\omega}$ in figure 2, the results in figure 18 hint that polymers act to break the mirror symmetry, and that this tendency increases with their elasticity.
7 Conclusion
In this work, we have investigated the cascades of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence. The study is based on direct numerical simulations of a dilute polymeric solution at $Re_{\lambda}\approx 460$, with $1/9\leq De\leq 9$. We have extended the formulation introduced by Baj et al. (2022) and derived the exact scale-by-scale budget equations for the energy, enstrophy and helicity, $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, for polymeric turbulent flows. The equations are general and remain valid for a generic inhomogeneous and anisotropic turbulent flow as well. They capture the mechanisms of production, transfer and dissipation in the combined space of scales and positions. Compared to purely Newtonian flows, the polymer/fluid interaction introduces additional sink/source process and an alternative scale-space transfer for the three quantities. The newly derived equations have been then tailored to homogeneous and isotropic turbulence for the present problem.
For $De\geq 1$ polymers effectively modify the cascades of energy and helicity, the two inviscid invariants of the $3D$ Navier–Stokes equations. Both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ are injected in the system at the largest scales and transferred to the small scales by two distinct transfer processes driven by fluid inertia and fluid/polymer interactions, where they are dissipated away. At large scales, the nonlinear cascade dominates, while at small scales the polymer-driven transfer takes over. The cross-over scale between the two dominant processes agrees fairly well $r_{p}^{*}$ at which $\tau_{p}/\tau_{f}(r_{p}^{*})=1$, where $\tau_{p}$ and $\tau_{f}(r)$ are the polymer relaxation time and the characteristic turnover time of eddies at scale $r$. When $r<r_{p}^{*}$ and $\tau_{p}>\tau_{f}(r)$ polymers effectively interact with the $r$-fluctuations of the carrier flow. The coexistence of the two transfer mechanisms leads to a multiscaling behaviour of the velocity and helicity structure functions. At small $De$ and large $r$, $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ exhibit the same power law predicted by the Kolmogorov theory, i.e. $\left\langle{\delta q^{2}}\right\rangle\sim\left\langle{\delta h}\right\rangle\sim r^{2/3}$ with $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle\approx 0.75$, consistently with helicity behaving as a passive scalar in HIT (Borue & Orszag, 1997). For $De\geq 1$ and at smaller scales, instead, both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ deviate from the Kolmogorov predictions, with the latter showing a steeper slope. For $De=1$ we measure $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ (in agreement with Zhang et al., 2021) and $\left\langle{\delta h}\right\rangle\sim r^{0.9}$. Accordingly, we observe that compared to the purely Newtonian case, the relative helicity $r\left\langle{\delta h}\right\rangle/2\left\langle{\delta q^{2}}\right\rangle$ increases with $De$ at all scales, indicating that polymers favour events that break mirror symmetry at small scales.
A closer look at the energy fluxes reveals that polymers deplete the nonlinear cascade by weakening both the direct and inverse extreme events of both cascades, leading to a less skewed flux distribution as $De$ increases. For $r\gtrapprox r_{p}^{*}$, the nonlinear flux distribution collapses reasonably well for all $De$ once the quantities are normalised with their standard deviation. Moving to the polymeric flux, we observe that the amount of energy carried by the polymer-driven cascade increases with $De$ at all scales. Similar to the nonlinear flux, on average the polymeric flux transfers energy from larger to smaller scales, but local events with intense backscatter exist which transfer energy from small to large scales. However, the probability of inverse transfer decreases as $De$ increases.
Unlike energy and helicity, enstrophy is not an inviscid invariant of the $3D$ Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, which is related to the enstrophy produced by vortex stretching-like processes (Davidson, 2004). The conventional notion of cascade used for energy and helicity, thus, does not apply in this case, as most of the enstrophy is directly generated at the small scales. However, vortex stretching is active in the inertial range which results in an increasing rate of transfer of $\left\langle{\delta\omega^{2}}\right\rangle$ at small $r$, being maximum at the end of the inertial range. For $De\geq 1$, the transfer via the polymeric route is comparable with the fluid nonlinear contribution and, unlike for $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, there is no transition with the two mechanisms coexisting in the entire inertial range. Overall, the net effect of polymers is to largely weaken vortex stretching. At large $De$, particularly, viscous dissipation balances the enstrophy generation due to fluid/polymer interaction. The underlying local flow topology further reveals that vortex stretching modulation is a result of polymers promoting events with a two-dimensional state of straining, like shear and planar extensional flows. Accordingly, we observe that polymers favour events where rotation and stretching are equally strong (being typical of vortex sheets and shear layers), rather than those dominated by rigid rotation.
Having characterised the influence of polymer additives on the simultaneous transfer of energy, enstrophy and helicity in polymeric homogeneous isotropic turbulence, the present study will serve as a stepping stone for similar investigations in more complex settings that may serve to elucidate the underlying mechanism of polymeric drag reduction. With a look towards applications, the next step is to introduce inhomogeneity and/or anisotropy in the flow, and use the presented formulation to investigate the influence of the polymers on the transfers of $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ in the combined space of scales and positions. In order of complexity, we mention for example shear polymeric turbulence (Robert et al., 2010; Warwaruk & Ghaemi, 2024), polymer-laden turbulent channels (Min et al., 2003; Izbassarov et al., 2021; Foggi Rota et al., 2024), and jets (Guimarães et al., 2020; Soligo & Rosti, 2023).
Acknowledgments
The authors acknowledge the computer time provided by the Scientific Computing and Data Analysis section of Research Support Division at OIST, and by HPCI, under the Research Project grants hp210269, hp220099, and hp230018.
Funding
The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding to M.E.R. from the Cabinet Office, Government of Japan. M.E.R. acknowledges funding from the Japan Society for the Promotion of Science (JSPS), grant 24K17210 and 24K00810.
Declaration of Interests
The authors report no conflict of interest.